### v3

# Reciprocal Beyond-Diagonal Reconfigurable Intelligent Surface  (BD-RIS):  Scattering Matrix Design via Manifold Optimization

Abstract
Beyond-diagonal reconfigurable intelligent surfaces (BD-RISs) are emerging as a transformative technology in wireless communications, enabling enhanced performance and quality of service (QoS) of wireless systems in harsh urban environments due to their relatively low cost and advanced signal processing capabilities. Generally, BD-RIS systems are employed to improve robustness, increase achievable rates, and enhance energy efficiency of wireless systems in both direct and indirect ways. The direct way is to produce a favorable propagation environment via the design of optimized scattering matrices, while the indirect way is to reap additional improvements via the design of multiple-input multiple-output (MIMO) beamformers that further exploit the latter “engineered” medium. In this article, the problem of sum-rate maximization via BD-RIS is examined, with a focus on feasibility, namely low-complexity physical implementation, by enforcing reciprocity in the BD-RIS design. We begin by outlining the system model and formulating an optimization problem that aims to enhance the system’s sum-rate by designing a symmetric scattering matrix. In particular, the approach leverages a manifold optimization framework, where a penalty term is added to the objective function to ensure that the symmetry constraint is upheld, with reciprocity further enforced by projecting the obtained solution onto a set of feasible scattering matrices. Simulation results demonstrate the effectiveness of the proposed method in outperforming current state-of-the-art (SotA) approaches in terms of sum-rate maximization.
I Introduction
Intelligent metasurfaces, or reconfigurable intelligent surfaces (RISs), are an emerging resource in the area of wireless communications, offering a new paradigm for controlling foundational features of electromagnetic fields to enhance system performance [1, 2, 3, 4, 5, 6, 7, 8]. These surfaces consist of individual elements, $i.e.$, passive circuits, with controllable properties that allow them to filter and reradiate the incoming signals without amplification, which is accomplished by implementing delays, reducing the amplitude, or changing wave polarization via the circuit’s local reprogrammable impedance.
Thanks to their low-power and reconfigurable nature, the so-called wave-domain signal processing approach offered by RISs offer several key advantages over classical analog/digital signal processing, including ease of deployment and substantial reductions in end-to-end latency and processing delay [9]. Motivated by these advantages, RISs have attracted enormous attention as an effective solution for enhancing wireless channel quality and consequently the communication performance of multiple-input single-output (MISO), multiple-input multiple-output (MIMO), and integrated sensing and communications (ISAC) systems [1]. Indeed, intelligent surfaces revolutionize wireless systems design by allowing dynamic reconfiguring of the propagation environment, achieving substantial gains without the need for additional active radio chains [10, 11, 12, 13].
Early research on the approach predominantly focused on diagonal reconfigurable intelligent surface (D-RIS) architectures, whose scattering matrices enabled simpler analysis and straightforward implementation [14, 15, 16, 17, 18, 19, 20, 21, 22]. Their use has been extensively studied across various communication settings under various design objectives, including sum-rate maximization [14], spectral efficiency improvement [15], energy efficiency enhancement [16], transmit power reduction [17], [18], and physical layer security [19]. To mention a few illustrative examples, in [20] a weighted sum-rate maximization problem under perfect and imperfect channel state information (CSI) was proposed, with algorithms developed for each case. Alternatively, in [21], an active RIS-aided downlink rate-splitting multiple access system is proposed, and its sum-rate performance was analyzed via a novel algorithm. To this extent, a joint optimization problem was formulated which aimed at designing the transmit (TX) beamforming and rate allocation vectors together with the RIS scattering matrix. The approach involves a two-stage design, where fractional programming (FP) techniques were used in combination with quadratic constraint quadratic programming (QCQP) methods to optimize the respective variables. Another excellent example is [19], whereby the authors utilize the D-RIS in an ISAC framework, applying manifold optimization to jointly improve sum-rate performance and enhance security against eavesdropping.
Although the conventional diagonal architecture offers analytical convenience, it also imposes structural constraints that limit the degrees-of-freedom (DoF), restricting its ability to manipulate the signal space. To overcome these limitations, subsequent studies have focused on beyond-diagonal reconfigurable intelligent surface (BD-RIS) architectures, which significantly increase the DoF available for signal shaping [23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33]. Consequently, manifold optimization and fractional programming have been leveraged as powerful tools to improve overall performance gains.
Junior et al. [23] provide a comprehensive tutorial on manifold optimization techniques and their application in RIS-aided massive MIMO (mMIMO) systems. They emphasize that exploiting the geometry of the constraint set typically yields more compelling solutions compared to standard methods of optimization, motivating the approach adopted in this article. Several recent works have applied these principles to specific BD-RIS design problems, employing manifold optimization to enhance secrecy rate [24] and to jointly configure the scattering matrix and TX precoder [25]. Both studies, however, are conducted in a non-reciprocal setting, $i.e.$, without enforcing symmetry.
One strategy for enforcing reciprocity via unitary and symmetry constraints implemented through auxiliary variables and objective penalties is presented in [28]. In that contribution, however, the reciprocity enforcement is somewhat cumbersome, as the constraints are imposed through a decoupling assisted by auxiliary variables, potentially leading to suboptimal results. In contrast, in the approach taken in this article, symmetry is ensured via a direct manifold optimization formulation of the sum-rate maximization problem.
Alternative BD-RIS design strategies extend beyond direct optimization from the sum-rate objective and instead focus on hybrid approaches. For example, in [26] two methods for designing BD-RIS scattering matrices are proposed. The first method seeks to maximize the total signal power received by the users following a passive maximum ratio transmission (MRT) strategy, while the second approach aims at nullifying the multi-user interference via a passive zero forcing (ZF) strategy. In contrast, this work adopts a fundamentally different strategy: rather than decoupling the design into separate MRT or ZF criteria, the scattering matrix is optimized with respect to the sum-rate objective. This direct formulation accounts for both signal enhancement and interference suppression, enabling these aspects to be addressed simultaneously within a unified and coherent optimization framework.
Another notable example is presented by Fang et al. [27], where the authors propose an alternative methodology—leveraging convex relaxation techniques—to address the non-convex BD-RIS problem. Specifically, they reformulate the non-convex constraints into a convex set, enabling efficient solution of the relaxed problem. The solution is then mapped back to the feasible region of the original non-convex set using their proposed symmetric unitary projection.
More recently, research has shifted toward non-reciprocal BD-RIS (NRBD-RIS) design within a full-duplex framework [30, 31], with Liu et al. [30] investigating their application in wireless circulator systems. Their results indicate that properly optimized non-reciprocal scattering matrices can yield improvements in both spectral efficiency and interference management compared to existing reciprocal structures, further expanding the applicability of BD-RIS in wireless networks.
Although prior studies, such as [25, 34, 35], incorporate manifold optimization into the scattering matrix design, to the best of our knowledge, no previous work simultaneously addresses the symmetry and unitary constraint of the BD-RIS within a manifold optimization framework directly applied to the sum-rate objective. This novel approach forms the basis for our proposed methodology, which will be discussed in detail in the following sections. All in all, the main contributions of this article are summarized as follows:
  • A novel manifold optimization method with symmetry enforcement is proposed to design the scattering matrix of BD-RIS. The proposed method, dubbed reciprocal BD-RIS (RBD-RIS), is based on a conjugate gradient ascent (CGA) algorithm, and aims to maximize the sum-rate of a BD-RIS-aided multi-user (MU)-MISO system.
  • The closed-form expression for the gradient of the objective function of the aforementioned problem is provided.
  • Thanks to the above, a generalized closed-form solution of the RBD-RIS design problem is given, which can be particularized for group-connected cases of arbitrary sizes, including single- and fully-connected architectures.
  • Simulation results showing that the proposed method outperforms state-of-the-art (SotA) counterparts [26] are provided, which demonstrate that reciprocity in the scattering matrices does not implicate in performance loss when parameterizing BD-RISs.
Organization: Section II introduces the BD-RIS-aided MU-MISO system model, presents the different BD-RIS architectures considered, and formulates the problem considered in this article. Next, a general solution to the first stage of the optimization problem, namely the scattering matrix design is provided in Section III. Simulation results are offerd in Section IV, which help evaluate the performance of the proposed designs. Finally, a brief conclusion is given in Section V, followed by Appendix A and B, where proofs of two Lemmas are detailed.
Notation: Unless otherwise specified, $\mathbf{X}$ and $\mathbf{x}$ denote matrices and vectors. The absolute value, $\ell^{2}$ and Frobenius norm are denoted by $|\cdot|$, $\|\cdot\|_{2}$, $\|\cdot\|_{F}$. The transpose and the Hermitian transpose are denoted by $(\cdot)^{T}$ and $(\cdot)^{H}$. In addition, $\mathbf{X}_{i,j}$ denotes the $i$-th row and $j$-th column element of the matrix $\mathbf{X}$, $\mathbf{X}_{i:\bar{i},j:\bar{j}}$ extracts the elements from the $i$-th to the $\bar{i}$-th row, and $j$-th to $\bar{j}$ column from the matrix $\mathbf{X}$. Furthermore, $\text{diag}(\mathbf{x})=(x_{1},x_{2},\dots,x_{k})$ denotes a diagonal matrix, where the main diagonal is $\mathbf{x}$, and $\text{blkdiag}\left(\mathbf{X}_{1},\mathbf{X}_{2},\dots,\mathbf{X}_{k}\right)$ denotes a $G\times G$ block diagonal matrix with off diagonal elements 0, where the blocks are $\mathbf{X}_{1},\mathbf{X}_{2},\dots,\mathbf{X}_{k}$. The sets of complex and real numbers are denoted by $\mathbb{C}$ and $\mathbb{R}$, while the real and imaginary values of a complex number are denoted by $\Re\{\cdot\}$ and $\Im\{\cdot\}$, respectively. The circular-symmetric complex normal random variable with $\sigma^{2}$ variance is denoted as $\mathcal{CN}(0,\sigma^{2})$.
II System Model
Consider a BD-RIS aided downlink MU-MISO system as illustrated in Figure 1, where a base station (BS) with $N$ TX antennas serves $K$ single receive (RX) antenna users [^1] This scenario is assumed in accordance to SotA counterparts., denoted as $U_{k}$, $\forall k\in\{1,2,\dots,K\}$, and a BD-RIS consisting of $R$ reflective elements (REs). Hence, the BD-RIS scattering matrix is represented by $\mathbf{\Theta}\in\mathbb{C}^{R\times R}$. The channel linking the BS and the BD-RIS ($i.e.$, the BS-BD-RIS channel) is denoted as $\mathbf{H}_{\mathrm{TX}}\in\mathbb{C}^{R\times N}$, and the BD-RIS-$U_{k}$ channel vector is $\mathbf{h}_{k}\in\mathbb{C}^{R\times 1}$.
Figure 1: Illustration of the system model, where a BS with $N$ TX antennas serves $K$ single-antenna users through an $R$-Element RBD-RIS, without a LoS link2 between the BS and the users.
The transmit signal vector is given by $\mathbf{x}=\mathbf{V}\mathbf{s}$, where the information symbols $\mathbf{s}\in\mathbb{C}^{K\times 1}$ satisfy $\mathbb{E}[\mathbf{s}\mathbf{s}^{H}]=\mathbf{I}$, and the beamforming matrix $\mathbf{V}\in\mathbb{C}^{N\times K}$ meets the power constraint $\|\mathbf{V}\|_{F}^{2}\leq P_{\mathrm{max}}$, with $P_{\mathrm{max}}$ being the maximum transmit power at the BS. It is assumed that the BS-$U_{k}$ channel is blocked [^2] This assumption is made only to enable direct comparison of the proposed technique with SotA methods, $e.g.$ [26], and is not required by our approach, since the scattering matrix design depends only on the cascaded BS-RIS-to-user channels. However, we are aware that overall performance under such assumptions relies on the BS-RIS channel having a sufficiently high rank, which holds only in rich-scattering, large-antenna setups, motivating an extension of the work under poorly scattered channels with LoS links between the BS and users. Such an extension also requires, however, further work on the beamforming algorithm, thus falling beyond the scope of this article, and therefore will be pursued in a follow-up contribution., such that a focus is placed on the performance gains due to the BD-RIS. Thus, the signal $r_{k}\in\mathbb{C}$ received at $U_{k}$ can be expressed as
$r_{k}=\mathbf{h}_{k}^{T}\mathbf{\Theta}\mathbf{H}_{\mathrm{TX}}\mathbf{x}+n_{k},$ (1)
where $n_{k}\sim\mathcal{CN}(0,N_{0})$ denotes additive white Gaussian noise (AWGN) with power $N_{0}$.
The compact vector representation in (1) is inspired by the formulation used in [26], and is given as
$\mathbf{r}=[r_{1},r_{2},\dots,r_{K}]^{T},$ (2a)
$\mathbf{n}=[n_{1},n_{2},\dots,n_{K}]^{T},$ (2b)
$\mathbf{V}=[\mathbf{v}_{1},\mathbf{v}_{2},\dots,\mathbf{v}_{K}],$ (2c)
$\mathbf{H}_{\mathrm{RX}}=[\mathbf{h}_{1},\mathbf{h}_{2},\dots,\mathbf{h}_{K}]^{T},$ (2d)
$\mathbf{H}_{\mathrm{TX}}=\left[\mathbf{w}_{1},\mathbf{w}_{2},\dots,\mathbf{w}_{N}\right],$ (2e)
$\mathbf{r}=\mathbf{H}_{\mathrm{RX}}\boldsymbol{\Theta}\mathbf{H}_{\mathrm{TX}}\mathbf{x}+\mathbf{n},$ (2f)
where $\mathbf{r}\in\mathbb{C}^{K\times 1}$, $\mathbf{H}_{\mathrm{RX}}\in\mathbb{C}^{K\times R}$, and $\mathbf{w}_{n}\in\mathbb{C}^{R\times 1}$, s.t. $n\in\{1,2,\dots,N\}$, denote the received signal vector, BD-RIS-$U_{k}$ channel matrix, and BS-BD-RIS channel vector, respectively.
II-A Scattering Matrix Definition
Depending on the reconfigurable impedance network of the BD-RIS, three different architectures are considered, namely:
  1. 1.
    Single-connected BD-RIS: Equivalent to the conventional D-RIS, where each RE is independently connected to a single grounded reconfigurable impedance with no interconnections among REs. Consequently, the resulting scattering matrix is inherently diagonal, which naturally satisfies the symmetry constraints, while also being unitary due to the reciprocal and lossless nature. The aforementioned constraints are described as
    $\mathcal{S}_{\text{SC}_{1}}=\left\{\mathbf{\Theta}:[\mathbf{\Theta}]_{i,j}=0,\ \forall i\neq j\right\},$ (3)
    $\mathcal{S}_{\text{SC}_{2}}=\left\{\mathbf{\Theta}:\left|[\mathbf{\Theta}]_{i,j}\right|=1,\ \forall i=j\right\},$ (4)
    where $i,j\in\{1,2,\dots,R\}$.
  2. 2.
    Fully-connected BD-RIS: In this architecture, every RE is connected to all others via a reconfigurable impedance network, allowing a wave incident on one element to be reflected by all others. As a result, the corresponding scattering matrix is full and—similarly to the other architectures—must satisfy both the symmetry and the unitary constraint, which for this case are given as
    $\mathcal{S}_{\text{FC}_{1}}=\left\{\mathbf{\Theta}:\mathbf{\Theta}=\mathbf{\Theta}^{T}\right\},$ (5)
    $\mathcal{S}_{\text{FC}_{2}}=\left\{\mathbf{\Theta}:\mathbf{\Theta}\mathbf{\Theta}^{H}=\mathbf{I}\right\}.$ (6)
  3. 3.
    Group-connected BD-RIS: In this intermediate architecture the REs are partitioned into $G$ disjoint groups, each consisting of $R_{G}=\frac{R}{G}$ elements. Within each group, all REs are interconnected, enabling waves impinging on one element to be reflected by all others in the same group. Apart from the intra-group connections, no other link exists. Consequently, the scattering matrix can be represented as a block-diagonal matrix by ordering its elements group-by-group, such that each block corresponds to one group and must independently satisfy the unitary and symmetry constraints. Notably, the group-connected architecture generalizes the previous two cases, reducing to the fully-connected case when $G=1$ and to the single-connected case when $G=R$. Denoting the scattering matrix of groups $g\in\{1,2,\ldots,G\}$ by $\mathbf{\Theta}_{g}\in\mathbb{C}^{R_{G}\times R_{G}}$, the constraints on the group-connected BD-RIS scattering matrix can be described as
    $\hskip-21.52771pt\mathcal{S}_{\text{GC}_{1}}\!\!=\!\left\{\mathbf{\Theta}\!:\!\mathbf{\Theta}\!=\!\mathrm{blkdiag}(\mathbf{\Theta}_{1},\ldots,\mathbf{\Theta}_{G}),\!\mathbf{\Theta}_{g}\!=\!\mathbf{\Theta}_{g}^{T},\forall g\right\}\!,\!\!\!$ (7)
    $\mathcal{S}_{\text{GC}_{2}}=\left\{\mathbf{\Theta}:\mathbf{\Theta}_{g}\mathbf{\Theta}_{g}^{H}=\mathbf{I}_{R_{G}},\;\forall g\right\}.$ (8)
II-B Problem Formulation
Given $\mathbf{\Theta}$ and $\mathbf{V}$, and assuming perfect channel state information, an achievable rate at $U_{k}$ is $\log_{2}(1+\gamma_{k})$, where the signal to interference-plus-noise ratio (SINR) is expressed as
$\gamma_{k}=\frac{\left|\mathbf{h}_{k}^{T}\mathbf{\Theta}\mathbf{H}_{\mathrm{TX}}\mathbf{v}_{k}\right|^{2}}{\sum\limits_{i\neq k}\left|\mathbf{h}_{k}^{T}\mathbf{\Theta}\mathbf{H}_{\mathrm{TX}}\mathbf{v}_{i}\right|^{2}+N_{0}}\ .\vskip-4.30554pt$ (9)
The corresponding sum-rate maximization problem is then
$(\text{P1}):\quad\underset{\mathbf{V},\mathbf{\Theta}}{\mathrm{maximize}}\quad$ $\sum_{k}\log_{2}(1+\gamma_{k})$ (10a)
subject to $\|\mathbf{V}\|_{F}^{2}\leq P_{\mathrm{max}},$ (10b)
$\mathbf{\Theta}\in\mathcal{S}_{a_{1}},$ (10c)
$\mathbf{\Theta}\in\mathcal{S}_{a_{2}},$ (10d)
where $a\in\{\text{SC},\text{FC},\text{GC}\}$, as described in [26].
The optimization problem (P1) is inherently non-convex due to the objective function in (10a), and the strong coupling between the TX beamforming and the scattering matrix. Additionally, the constraints in (10b)-(10d) are generally non-convex. As previously mentioned, the constraints in (10c) and (10d) reflect the BD-RIS’s reciprocal and lossless properties, respectively, with the constraint in (10d) representing the Stiefel manifold condition.
Due to this coupling, a two-stage design is adopted to enhance the sum-rate. In this article, the first stage of the optimization problem is addressed, namely, the design of the scattering matrix. The subsequent stage, focusing on beamforming matrix design, will be presented in a follow-up work.
III Stage 1: Scattering Matrix Design
We will solve (P1) in an iterative manner. To first design the scattering matrix, the impact of the BS beamforming is fixed by initializing it with a minimum mean square error (MMSE)-based beamformer. Accordingly, (2f) is rewritten as
$\mathbf{r}=\mathbf{H}_{\mathrm{RX}}\mathbf{\Theta}\mathbf{H}_{\mathrm{TX}}\mathbf{V}\mathbf{s}+\mathbf{n}\triangleq\mathbf{H}_{\mathrm{RX}}\mathbf{\Omega}\mathbf{V}\mathbf{s}+\mathbf{n}\triangleq\mathbf{E}\mathbf{V}\mathbf{s}+\mathbf{n},$ (11)
where $\mathbf{\Omega}=\mathbf{\Theta}\mathbf{H}_{\mathrm{TX}}=[\boldsymbol{\omega}_{1},\dots,\boldsymbol{\omega}_{N}]\in\mathbb{C}^{R\times N}$ and $\mathbf{E}=\mathbf{H}_{\mathrm{RX}}\mathbf{\Omega}=[\mathbf{e}_{1},\dots,\mathbf{e}_{K}]^{T}\in\mathbb{C}^{K\times N}$ denotes the equivalent channel. As a consequence, the SINR at $U_{k}$ during the design of the scattering matrix is given as
$\hskip-4.30554pt\gamma_{k}=\frac{\left|\mathbf{h}_{k}^{T}\boldsymbol{\Omega}\mathbf{v}_{k}\right|^{2}}{\sum\limits_{i\neq k}\left|\mathbf{h}_{k}^{T}\boldsymbol{\Omega}\mathbf{v}_{i}\right|^{2}+N_{0}}=\frac{\left|\mathbf{e}_{k}\mathbf{v}_{k}\right|^{2}}{\sum\limits_{i\neq k}\left|\mathbf{e}_{k}\mathbf{v}_{i}\right|^{2}+N_{0}}.$ (12)
Making use of (12), the optimization problem is formulated as
$(\text{P2}):\quad\underset{\mathbf{\Theta}}{\mathrm{maximize}}\quad$ $\sum_{k}\log_{2}(1+\gamma_{k})$ (13a)
subject to $\mathbf{\Theta}\in\mathcal{S}_{a_{1}},$ (13b)
$\mathbf{\Theta}\in\mathcal{S}_{a_{2}}.$ (13c)
III-A Solution for the Group-Connected architecture
As the group-connected BD-RIS architecture is the most generalized version of the three architectures considered, it can provide a generic solution to (P2), which in turn can be rewritten as
$(\text{P2a}):\quad\underset{\mathbf{\Theta}}{\mathrm{maximize}}\quad$ $\sum_{k}\log_{2}(1+\gamma_{k})$ (14a)
subject to $\mathbf{\Theta}_{g}=\mathbf{\Theta}_{g}^{T},$ (14b)
$\mathbf{\Theta}_{g}\mathbf{\Theta}_{g}^{H}=\mathbf{I}_{R_{G}},$ (14c)
where $g=\{1,2,\dots,G\}$. For simplicity, the sum-rate over all users is denoted as
$\eta=\sum_{k}\eta_{k},\quad\text{with}\quad\eta_{k}=\log_{2}(1+\gamma_{k}).$ (15)
As previously noted, the optimization problem (P2a) is inherently non-convex due to the SINR being a ratio of quadratic terms of $\mathbf{\Theta}$. The unitary equality constraint further complicates the problem due to its non-convex nature.
To address this matter, manifold optimization is employed by treating $\mathbf{\Theta}_{g}$ as a point on the Stiefel manifold. Meanwhile, the symmetry constraint restricts $\mathbf{\Theta}_{g}$ to the manifold of symmetric complex matrices, which forms a linear subspace and thus also a smooth manifold. To simultaneously enforce both the unitary and symmetry constraints, an intersection of the two sets is needed. However, as this intersection is challenging to handle directly, a workaround approach is considered by modifying the objective function to add a penalty that promotes symmetry.
Accordingly, the optimization problem then becomes
$\hskip-5.8125pt(\text{P3}):\quad\underset{\mathbf{\Theta}}{\mathrm{maximize}}\quad$ $\!\!\sum_{k}\log_{2}(1+\gamma_{k})-\nu\bigl\|\boldsymbol{\Theta}-\boldsymbol{\Theta}^{T}\bigr\|_{F}^{2}$ (16a)
subject to $\!\!\mathbf{\Theta}_{g}\mathbf{\Theta}_{g}^{H}=\mathbf{I}_{R_{G}},$ (16b)
where $\nu\in\mathbb{R}$ denotes a nonnegative weight and $\bigl\|\boldsymbol{\Theta}-\boldsymbol{\Theta}^{T}\bigr\|_{F}^{2}$ is used to obtain a quantitative measure of how far the matrix is from being symmetric.
Having defined the optimization problem (P3), its solution can be obtained numerically using the Manopt toolbox [36]. However, to reduce the high computational complexity associated with Manopt, the corresponding CGA algorithm [37], tailored for this problem, is derived in the sequel. This algorithm incorporates a closed-form expression for the gradient of the objective function, as seen in (16a), enabling a more efficient implementation of the scattering matrix design.
To this extent, the group-wise definition of the system model is considered, namely
$\mathbf{\Theta}_{g}=\mathbf{\Theta}_{[R_{G}(g-1)+1\,:\,gR_{G},\;\;R_{G}(g-1)+1\,:\,gR_{G}]},$ (17)
where $g\in\{1,2,\dots,G\}$.
As such, the group-wise SINR  [^3] This notation considers only the contributions of $\mathbf{\Theta}_{g}$, with the contributions from all other groups set to zero. equation can be written as
$\gamma_{k}^{(g)}=\frac{\left|\mathbf{h}_{k}^{(g)T}\mathbf{\Theta}_{g}\mathbf{W}^{(g)}\mathbf{v}_{k}\right|^{2}}{\sum\limits_{i\neq k}\left|\mathbf{h}_{k}^{(g)T}\mathbf{\Theta}_{g}\mathbf{W}^{(g)}\mathbf{v}_{i}\right|^{2}+N_{0}},$ (18)
where $\mathbf{v}_{l}\in\mathbb{C}^{N\times 1}$ denotes the beamforming vector, with $l\in\{i,k\}$.
The vector $\mathbf{h}_{k}^{(g)T}=\mathbf{H}_{RX\left[k,R_{G}(g-1)+1:gR_{G}\right]}$ and the matrix $\mathbf{W}^{(g)}=\mathbf{H}_{TX\left[R_{G}(g-1)+1:gR_{G},1:N\right]}$ represent the channel components associated with the $g$-th group. Specifically, $\mathbf{h}_{k}^{(g)T}\in\mathbb{C}^{1\times R_{G}}$ denotes the BD-RIS-$U_{k}$ channel vector, and $\mathbf{W}^{(g)}\in\mathbb{C}^{R_{G}\times N}$ denotes the BS-BD-RIS channel matrix corresponding to the group size $R_{G}$, respectively. Within the $g$-th group dimension, the scattering matrix $\mathbf{\Theta}_{g}\in\mathbb{C}^{R_{G}\times R_{G}}$ models the group-specific scattering behavior. Accordingly, the group-wise equivalent channel is defined as $\mathbf{e}_{k}^{(g)}=\mathbf{h}_{k}^{(g)T}\mathbf{\Theta}_{g}\mathbf{W}^{(g)}$, where $\mathbf{e}_{k}^{(g)}\in\mathbb{C}^{1\times N}$, enabling the SINR to be rewritten as
$\gamma_{k}^{(g)}=\frac{\left|\mathbf{e}_{k}^{(g)}\mathbf{v}_{k}\right|^{2}}{\sum\limits_{i\neq k}\left|\mathbf{e}_{k}^{(g)}\mathbf{v}_{i}\right|^{2}+N_{0}}.$ (19)
The gradient of the objective function (16a) with respect to each group $\mathbf{\Theta}_{g}$ is computed as
$\!\!\!\nabla_{\mathbf{\Theta}_{g}}\!\eta$ $=\nabla_{\mathbf{\Theta}_{g}}\!\!\left(\!\sum_{k}\log_{2}(1+\gamma_{k}^{(g)})-\nu\left\|\mathbf{\Theta}_{g}-\mathbf{\Theta}_{g}^{T}\right\|_{F}^{2}\!\right)$ (20)
$=\frac{1}{\ln(2)}\sum_{k}\!\frac{1}{1+\gamma_{k}^{(g)}}\nabla_{\mathbf{\Theta}_{g}}\gamma_{k}^{(g)}\!-\nu\nabla_{\mathbf{\Theta}_{g}}\!\!\left\|\mathbf{\Theta}_{g}\!-\mathbf{\Theta}_{g}^{T}\right\|_{F}^{2}\!,$
where the symmetry penalty term can be rewritten as
$\left\|\mathbf{\Theta}_{g}-\mathbf{\Theta}_{g}^{T}\right\|_{F}^{2}$ $=\mathrm{Tr}\{(\mathbf{\Theta}_{g}-\mathbf{\Theta}_{g}^{T})^{H}(\mathbf{\Theta}_{g}-\mathbf{\Theta}_{g}^{T})\}$ (21)
$=\mathrm{Tr}\{(\mathbf{\Theta}_{g}^{H}-\mathbf{\Theta}_{g}^{\star})(\mathbf{\Theta}_{g}-\mathbf{\Theta}_{g}^{T})\}$
$=\mathrm{Tr}\{\mathbf{\Theta}_{g}^{H}\mathbf{\Theta}_{g}-\mathbf{\Theta}_{g}^{H}\mathbf{\Theta}_{g}^{T}-\mathbf{\Theta}_{g}^{\star}\mathbf{\Theta}_{g}+\mathbf{\Theta}_{g}^{\star}\mathbf{\Theta}_{g}^{T}\}.$
The gradient for half of the terms in (21) can be effectively found in [38], namely
$\nabla_{\mathbf{\Theta}_{g}}\mathrm{Tr}\{\mathbf{\Theta}_{g}^{H}\mathbf{\Theta}_{g}\}=\nabla_{\mathbf{\Theta}_{g}}\mathrm{Tr}\{\mathbf{\Theta}_{g}^{\star}\mathbf{\Theta}_{g}^{T}\}^{T}=2\mathbf{\Theta}_{g}.$ (22)
Conversely,
$\nabla_{\mathbf{\Theta}_{g}}\mathrm{Tr}\{\mathbf{\Theta}_{g}^{H}\mathbf{\Theta}_{g}^{T}\}=\nabla_{\mathbf{\Theta}_{g}}\mathrm{Tr}\{\mathbf{\Theta}_{g}^{\star}\mathbf{\Theta}_{g}\}^{T},$ (23)
where the gradient is obtained as
$\nabla_{\mathbf{\Theta}_{g}}\mathrm{Tr}\{\mathbf{\Theta}_{g}^{\star}\mathbf{\Theta}_{g}\}=2\frac{\partial\mathrm{Tr}\{\mathbf{\Theta}_{g}^{\star}\mathbf{\Theta}_{g}\}}{\partial\mathbf{\Theta}_{g}^{\star}}=2\mathbf{\Theta}_{g}^{T},$ (24)
using the result from Lemma 1.
The reader is referred to Appendix A for further details.
Based on the above, the gradient of $\left\|\mathbf{\Theta}_{g}-\mathbf{\Theta}_{g}^{T}\right\|_{F}^{2}$ with respect to $\mathbf{\Theta}_{g}$ is derived as
$\nabla_{\mathbf{\Theta}_{g}}\left\|\mathbf{\Theta}_{g}-\mathbf{\Theta}_{g}^{T}\right\|_{F}^{2}=4\left(\mathbf{\Theta}_{g}-\mathbf{\Theta}_{g}^{T}\right).$ (25)
For simplicity, the numerator and denominator of $\gamma_{k}$ in (19) are denoted as $\chi_{k}$ and $\psi_{k}$, respectively, $i.e.$,
$\gamma_{k}^{(g)}=\frac{\chi_{k}}{\psi_{k}}=\frac{\left|\mathbf{e}_{k}^{(g)}\mathbf{v}_{k}\right|^{2}}{\sum\limits_{i\neq k}\left|\mathbf{e}_{k}^{(g)}\mathbf{v}_{i}\right|^{2}+N_{0}},$ (26)
such that the gradient of $\gamma_{k}$ can be expressed as
$\nabla_{\mathbf{\Theta}_{g}}\gamma_{k}^{(g)}=\frac{\psi_{k}\nabla_{\mathbf{\Theta}_{g}}\chi_{k}-\chi_{k}\nabla_{\mathbf{\Theta}_{g}}\psi_{k}}{\psi_{k}^{2}},$ (27)
where the gradients of the numerator and denominator are computed as follows:
$\nabla_{\mathbf{\Theta}_{g}}\chi_{k}=\nabla_{\mathbf{\Theta}_{g}}\left|\mathbf{e}_{k}^{(g)}\mathbf{v}_{k}\right|^{2}=2\,\Bigg((\mathbf{e}_{k}^{(g)}\mathbf{v}_{k})^{\star}\frac{\partial\mathbf{e}_{k}^{(g)}\mathbf{v}_{k}}{\partial\mathbf{\Theta}_{g}}\Bigg)^{\star},$ (28)
$\nabla_{\mathbf{\Theta}_{g}}\psi_{k}$ $=\nabla_{\mathbf{\Theta}_{g}}\!\left(\sum_{i\neq k}\left|\mathbf{e}_{k}^{(g)}\mathbf{v}_{i}\right|^{2}+N_{0}\right)$ (29)
$=2\,\Bigg(\sum_{i\neq k}(\mathbf{e}_{k}^{(g)}\mathbf{v}_{i})^{\star}\frac{\partial\mathbf{e}_{k}^{(g)}\mathbf{v}_{i}}{\partial\mathbf{\Theta}_{g}}\Bigg)^{\star},$
using the result obtained in Lemma 2. The reader is referred to Appendix B for further details.
Finally, the partial derivatives seen in (28) and (29) are computed as
$\hskip-5.38193pt\frac{\partial\mathbf{e}_{k}^{(g)}\mathbf{v}_{l}}{\partial\mathbf{\Theta}_{g}}$ $=\frac{\partial}{\partial\mathbf{\Theta}_{g}}\!\left(\mathbf{h}_{k}^{(g)T}\mathbf{\Theta}_{g}\mathbf{W}^{(g)}\mathbf{v}_{l}\right)\!=\mathbf{h}_{k}^{(g)}(\mathbf{W}^{(g)}\mathbf{v}_{l})^{T}\!.$ (30)
Accordingly, the gradient of the objective function in (20) is reformulated as shown in (31). For convenience, (32) presents the gradient of the objective function assuming a simple power allocation matrix used for beamforming [^4] This assumption can only be made for fully-loaded systems ($K$ = $N$)., $i.e.$, $\mathbf{V}=\mathrm{diag}(\sqrt{v}_{1},\dots,\sqrt{v}_{K})$, where $v_{k}\in\mathbb{R},\forall k$. Both gradients are provided at the top of the following page.
$\nabla_{\mathbf{\Theta}_{g}}\eta=\!\sum_{k}\frac{2\Big(\sum\limits_{i\neq k}\left|\mathbf{e}_{k}^{(g)}\mathbf{v}_{i}\right|^{2}\!\!+\!N_{0}\Big)\!\Big(\!(\mathbf{e}_{k}^{(g)}\mathbf{v}_{k})^{\star}\mathbf{h}_{k}^{(g)}(\mathbf{W}^{(g)}\mathbf{v}_{k})^{T}\Big)^{\star}\!\!-2\!\left|\mathbf{e}_{k}^{(g)}\mathbf{v}_{k}\right|^{2}\!\!\sum\limits_{i\neq k}\!\Big(\!(\mathbf{e}_{k}^{(g)}\mathbf{v}_{i})^{\star}\mathbf{h}_{k}^{(g)}(\mathbf{W}^{(g)}\mathbf{v}_{i})^{T}\Big)^{\star}}{\ln(2)(1\!+\!\gamma_{k}^{(g)})\Big(\sum\limits_{i\neq k}\left|\mathbf{e}_{k}^{(g)}\mathbf{v}_{i}\right|^{2}\!\!+N_{0}\Big)^{2}}\!-4\nu(\mathbf{\Theta}_{g}\!-\mathbf{\Theta}_{g}^{T})$ (31)
$\nabla_{\mathbf{\Theta}_{g}}\eta=\sum_{k}\frac{2v_{k}\Big(v_{k}\sum\limits_{i\neq k}\left|e_{k,i}^{(g)}\right|^{2}\!\!+N_{0}\Big)\Big(e_{k,k}^{(g)\star}\mathbf{h}_{k}^{(g)}\mathbf{w}_{k}^{(g)T}\Big)^{\star}\!\!-2v_{k}^{2}\left|e_{k,k}^{(g)}\right|^{2}\!\sum\limits_{i\neq k}\Big(e_{k,i}^{(g)\star}\mathbf{h}_{k}^{(g)}\mathbf{w}_{i}^{(g)T}\Big)^{\star}}{\ln(2)(1\!+\!\gamma_{k}^{(g)})\Big(v_{k}\sum\limits_{i\neq k}\left|e_{k,i}^{(g)}\right|^{2}\!\!+N_{0}\Big)^{2}}\!-4\nu(\mathbf{\Theta}_{g}\!-\mathbf{\Theta}_{g}^{T})$ (32)
The constraint of the optimization problem, given in (16b), defines a complex Stiefel manifold of dimension $R_{g}^{2}$. For clarity, a manifold is a topological space that locally resembles the Euclidean space, but may have a globally curved or more complex structure. Such manifolds naturally enforce constraints and symmetries, including orthogonality or unitary conditions [23].
The complex Stiefel manifold is defined as
$\mathrm{St}(n,p;\mathbb{C})=\left\{\mathbf{X}\in\mathbb{C}^{n\times p}\,:\,\mathbf{X}^{H}\mathbf{X}=\mathbf{I}_{p}\right\}.$ (33)
For square blocks ($p=n=R_{G}$), the Stiefel manifold $\mathrm{St}(R_{G},R_{G})$ is given by
$\mathrm{St}(R_{G},R_{G};\mathbb{C})=\{\mathbf{\Theta}_{g}\in\mathbb{C}^{R_{G}\times R_{G}}:\mathbf{\Theta}_{g}\mathbf{\Theta}_{g}^{H}=\mathbf{I}_{R_{G}}\}.$ (34)
Since the general group-connected BD-RIS scattering matrix is block-diagonal, it can be expressed as
$\hskip-1.89444pt\boldsymbol{\Theta}=\mathrm{blkdiag}(\boldsymbol{\Theta}_{1},\boldsymbol{\Theta}_{2},\dots,\boldsymbol{\Theta}_{G}),\>\boldsymbol{\Theta}_{g}\in\mathrm{St}(R_{G},R_{G};\mathbb{C}),$ (35)
where $g\in\{1,2,\dots,G\}$.
The usage of the CGA algorithm is motivated by the rich geometric structure of Riemannian manifolds, which allows well-defined gradients of cost functions. Additionally, optimization over a manifold is locally similar to that in the Euclidean space. As a result, optimization techniques developed for the Euclidean spaces, such as the gradient descent and trust-region methods, have corresponding formulations on manifolds [15].
The algorithm proceeds through a sequence of retraction operations $\mathrm{R}(\cdot,\cdot)$, followed by a projection $\mathrm{T}(\cdot,\cdot)$ onto the tangent space. The retraction operation is expressed as
$\mathrm{R}_{\mathbf{\Theta}}(\mathbf{\Theta},\boldsymbol{\Xi})=\mathcal{Q}(\mathbf{\Theta}+\alpha\boldsymbol{\Xi}),$ (36)
where $\alpha$ denotes the step size, $\mathbf{\Theta}$ the current point on the manifold, and $\boldsymbol{\Xi}$ the ascent direction in the tangent space.
In the above, the operator $\mathcal{Q}(\cdot)$ returns the Q-factor from the QR decomposition of $\mathbf{\Theta}+\alpha\boldsymbol{\Xi}$, ensuring that $\mathrm{R}_{\mathbf{\Theta}}(\mathbf{\Theta},\boldsymbol{\Xi})\in\mathrm{St}(R,R;\mathbb{C})$ [39]. Thus, the retraction function maps a point in the tangent space back onto the manifold.
The retraction function for each block is defined as
$\mathbf{Y}_{g}=\mathbf{\Theta}_{g}+\alpha\boldsymbol{\Xi}_{g},\>[\mathcal{Q}_{g},\mathcal{R}_{g}]=\mathcal{QR}(\mathbf{Y}_{g}),\>\mathbf{\Theta}_{g}^{\text{new}}=\mathcal{Q}_{g}.$ (37)
Complementary to the retraction operation, the tangent projection function describes the feasible directions of movement at a point on a manifold, represented by tangent vectors. All tangent vectors at a point form the tangent space, which includes every possible direction the point can move to.
Furthermore, the tangent space can be regarded as a Euclidean space, where a particular tangent vector, denoted as the Riemannian gradient, indicates the direction of steepest descent of the objection function [40].
In light of the above, it is initially required for the Euclidean gradient of the objective function to be computed. The Riemannian gradient is then obtained by projecting this Euclidean gradient, as given in (31) and (32), onto the tangent space. This projection is expressed as
$\mathrm{T}_{\mathbf{\Theta}}(\nabla_{\mathbf{\Theta}}\eta,\mathbf{\Theta})=\nabla_{\mathbf{\Theta}}\eta-\mathbf{\Theta}\cdot\frac{\mathbf{\Theta}^{H}\nabla_{\mathbf{\Theta}}\eta+(\nabla_{\mathbf{\Theta}}\eta)^{H}\mathbf{\Theta}}{2}.\vskip-2.15277pt$ (38)
Straightforwardly, for each group $g$, the projection onto the tangent space is given as
$\mathrm{T}_{\mathbf{\Theta}_{g}}(\nabla_{\!\mathbf{\Theta}_{g}}\eta,\mathbf{\Theta}_{g})\!=\!\nabla_{\!\mathbf{\Theta}_{g}}\eta-\mathbf{\Theta}_{g}\cdot\frac{\mathbf{\Theta}_{g}^{H}\nabla_{\mathbf{\Theta}_{g}}\eta+(\nabla_{\mathbf{\Theta}_{g}}\eta)^{H}\mathbf{\Theta}_{g}}{2}.\vskip-2.15277pt$ (39)
After the optimization, $i.e.$, following the sequence of retraction and tangent projection operations, the method from [27] is applied to enforce the symmetry and unitary constraints, since adding a penalty serves only as a workaround approach.
For the reader’s convenience, the projection method is summarized as follows. First, a symmetric matrix projection ensures that $\mathbf{A}=\mathbf{A}^{T}$, by averaging the solution with its transpose, $i.e.$,
$\Pi_{S_{sym}}(\mathbf{A})=\frac{1}{2}(\mathbf{A}+\mathbf{A}^{T}).\vskip-2.15277pt$ (40)
Second, to enforce the unitary constraint, the matrix is projected onto the set of unitary matrices using the singular value decomposition (SVD). Specifically, if $\mathbf{A}=\mathbf{U}\boldsymbol{\Sigma}\mathbf{V}^{H}$, the unitary projection is defined as
$\Pi_{S_{uni}}(\mathbf{A})=\mathbf{U}\mathbf{V}^{H}.\vskip-2.15277pt$ (41)
Lastly, when both symmetry and unitary constraints must be enforced simultaneously, the procedure is as follows. If $\mathbf{B}$ is defined as the result after the symmetric projection of $\mathbf{A}$, $i.e.$, $\mathbf{B}=\Pi_{S_{sym}}(\mathbf{A})$, then given that the SVD of $\mathbf{B}=\mathbf{U}\boldsymbol{\Sigma}\mathbf{V}^{H}$, the unitary symmetric projection of $\mathbf{A}$ is defined as
$\Pi_{S_{unisym}}(\mathbf{A})=\mathbf{U}\mathbf{V}^{H},$ (42)
where $\mathbf{U}$ and $\mathbf{V}^{H}$ denote the left and right singular vector matrices, respectively.
Algorithm 1 Proposed CGA for BD-RIS Optimization
[…]
0: Optimized RIS matrix $\mathbf{\Theta}^{\mathrm{opt}}$
[…]
1: Compute initial objective $\eta(\mathbf{\Theta}^{(0)})$ and Riemannian gradient $\mathbf{r}^{(0)}=\mathrm{T}_{\mathbf{\Theta}}(\nabla_{\mathbf{\Theta}}\eta(\mathbf{\Theta}^{(0)}),\mathbf{\Theta}^{(0)})$
2: Set initial search direction $\boldsymbol{\Xi}^{(0)}=-\mathbf{r}^{(0)}$
3:for $i=0$ to II do
4:  if $\langle\mathbf{r}^{(i)},\boldsymbol{\Xi}^{(i)}\rangle\leq 0$ then
5:   Set $\boldsymbol{\Xi}^{(i)}=\mathbf{r}^{(i)}$
6:  end if
7:  Compute $\alpha^{(i)}$ via Armijo line search and update $\mathbf{\Theta}^{(i+1)}$
8:  Compute new objective $\eta(\mathbf{\Theta}^{(i+1)})$, using (16a)
9:  Compute new Riemannian gradient $\mathbf{r}^{(i+1)}$, using (39)
10:  Compute $\beta^{(i)}=\max\left(0,\frac{\langle\mathbf{r}^{(i+1)},\mathbf{r}^{(i+1)}-\mathbf{r}^{(i)}\rangle}{\langle\mathbf{r}^{(i)},\boldsymbol{\Xi}^{(i)}\rangle}\right)$
11:  Update direction: $\boldsymbol{\Xi}^{(i+1)}=-\mathbf{r}^{(i+1)}+\beta^{(i)}\boldsymbol{\Xi}^{(i)}$
12:  if $|\eta(\mathbf{\Theta}^{(i+1)})-\eta(\mathbf{\Theta}^{(i)})|<\epsilon$ then
13:   break
14:  end if
15:end for
16:for $g=1$ to $G$ do
17:  Extract block $\mathbf{\Theta}_{g}$ from $\mathbf{\Theta}^{(i+1)}$
18:  Symmetrize: $\mathbf{\Theta}_{\mathrm{sym}}=\frac{1}{2}(\mathbf{\Theta}_{g}+\mathbf{\Theta}_{g}^{T})$, using (40)
19:  Perform SVD: $\mathbf{\Theta}_{\mathrm{sym}}=\mathbf{U}\mathbf{\Sigma}\mathbf{V}^{H}$, using (41)
20:  Set $\mathbf{\Theta}^{\mathrm{opt}}_{g}=\mathbf{U}\mathbf{V}^{H}$, using (42)
21:end for
22: Assemble full $\mathbf{\Theta}^{\mathrm{opt}}$ as block diagonal of all $\mathbf{\Theta}_{g}^{\mathrm{opt}}$
23:return $\mathbf{\Theta}_{\mathrm{opt}}$
The full procedure implemented in this article is outlined in Algorithm 1. The algorithm begins by initializing a random unitary symmetric block-diagonal matrix $\mathbf{\Theta}^{(0)}$ and computing the initial objective function along with its gradient. In each iteration, these quantities are updated according to the current values of the scattering matrix until convergence is achieved. The step size $\alpha$ is determined using an Armijo line search [40, 41], which ensures that there is a sufficient ascent. Furthermore, the parameter $\beta$ denotes the conjugate gradient (CG) momentum term, used to enhance the convergence speed of the algorithm.
Among various options, including the Fletcher-Reeves, Polak Ribière, and Hestenes-Stiefel formula, this algorithm incorporates the Polak-Ribière formula as described in [37, 41]. Finally, the scattering matrix resulting from the series of retraction and tangential projection operations is further processed using the method described in (40)-(42), ensuring that $\mathbf{\Theta}_{g}^{\mathrm{opt}}$ satisfies both the symmetric and unitary constraints.
Computational Complexity: The computational complexity of Algorithm 1 is primarily influenced by the Armijo line search and gradient computation. Within the line search, two operations dominate the complexity, namely a triple matrix multiplication and a QR decomposition-based retraction. Considering the general case of the group-connected architecture, the QR decomposition for each group, of size $R/G$, has a complexity of $\mathcal{O}((R/G)^{3})$, while the triple matrix multiplication has a complexity of $\mathcal{O}(KR(R+N))$. Since the QR decomposition is applied independently to each group, the total complexity of the retraction operation across all groups and users is given as $\mathcal{O}(KG(R^{3}/G^{3}))$.
Table I: Computational Complexity of BD-RIS Scattering Matrix Design
BD-RIS Architecture Computational Complexity
Interference Nulling SotA Alg. [26] Proposed Alg. 1
Single-connected $\mathcal{O}(tR+K^{6}+K^{4}R+K^{2}R^{2}+K^{2}R)$ $\mathcal{O}(IL(KR+K^{2}R^{2}+K^{2}RN))$
Group-connected $\mathcal{O}(tG\!R_{G}^{3}\!+\!tG\!R_{G}^{2}\!+\!K^{6}\!\!+\!K^{4}G\!R_{G}^{2}\!+\!K^{2}\!G^{2}\!R_{G}^{4}\!+\!K^{2}\!G\!R_{G}^{2})$ $\mathcal{O}(IL(KGR_{G}^{3}+K^{2}R^{2}+K^{2}RN))$
Fully-connected $\mathcal{O}(tR^{3}+tR^{2}+K^{6}+K^{4}R^{2}+K^{2}R^{4}+K^{2}R^{2})$ $\mathcal{O}(IL(KR^{3}+K^{2}R^{2}+K^{2}RN))$
In contrast, the triple matrix multiplication is performed on the full system, $i.e.$, not on a per-group basis, and yields a total complexity for all users of $\mathcal{O}(K(KR^{2}+KRN))$. With $L$ Armijo iterations, these contributions accumulate, giving an overall line search complexity of
$\mathcal{O}(LKG(R^{3}/G^{3})+LK(KR^{2}+KRN)).$ (43)
The gradient computation consists of several terms, resulting in a total complexity across all groups and users of
$\hskip-8.61108pt\mathcal{O}(KG(N^{2}\!+\!N\!R/G\!+\!(R/G)^{2}\!+\!K\!(R/G)^{2})\!+\!G(R/G)^{2}).\vskip-2.15277pt$ (44)
Furthermore, the initialization and projection described in (40)-(42) have a computational complexity of $\mathcal{O}(R^{3}/G^{2})$, which stems from the SVD computation. Lastly, the tangential projection in (39) has a complexity of $\mathcal{O}(R^{3})$. Considering $I$ iterations of the algorithm, and the case $L\gg 1$, the total complexity of the algorithm can then be expressed as
$\mathcal{O}(IL(KR^{3}/G^{2}+K^{2}R^{2}+K^{2}RN)).\vskip-2.15277pt$ (45)
Table I summarizes the simplified computational complexity associated with the different BD-RIS architectures, allowing for a straightforward comparison with SotA methods [^5] Notice that the SotA method in [26] assumes that the number of TX antennas at the BS equals the number of single-antenna users ($K=N$). Our method, does not require such assumption and therefore can be considered a generalization of the latter in this regard., where $t$ denotes the number of iterations required by the SotA algorithm [26]. The convergence behavior of Algorithm 1 is shown in Figure 2 for the different BD-RIS architectures considered, $i.e.$, single-, group-, with group size of 2 and 4, and fully-connected, respectively.
Figure 2: Convergence of Algorithm 1 vs. number of iterations II for different connectivity structures, with $P_{\mathrm{max}}=20\,\mathrm{dBm}$, $K=5$, $N=5$, and $R=32$.
For clarity, each BD-RIS architecture is represented by a distinct marker, as indicated in the legend of Figure 2, and this notation is used consistently throughout the article to facilitate cross-referencing of results.
The simulation settings are as follows: convergence tolerance of $\epsilon=10^{-8}$, maximum number of CGA iterations $I=8000$, maximum Armijo line search steps $L=200$, sufficient increase coefficient for the stepsize of $2\times 10^{-11}$, initial stepsize $c_{\text{init}}=1$, contraction factor for the stepsize $c_{\text{dec}}=0.75$, and the symmetry enforcing penalty constant $\nu=1$. Results indicate that more complex architectures require a larger number of iterations to converge; for instance, the single-connected case converges in approximately $50$ iterations, while the full-connected case requires around $2500$ iterations under the same conditions. Additionally, the convergence speed of the algorithm depends on the number of REs and the maximum transmit power $P_{\mathrm{max}}$, as well as on the initial scattering matrix, which in this article is chosen as a random unitary symmetric matrix. The figure represents results generated for one Monte-Carlo iteration, hence the reason for the non-smooth convergence curves.
As validated in Figure 2, the number $I$ of iterations required for convergence is considered for the scenario where $P_{\mathrm{max}}=20\mathrm{dBm}$, $K=5$ users, $N=5$ TX antennas, and $R=32$ REs. Specifically, these are $I_{\mathrm{SC}}=50$, $I_{\mathrm{GC}_{2}}=650$, $I_{\mathrm{GC}_{4}}=1050$ and $I_{\mathrm{FC}}=2500$, corresponding to the single-, group-, with group size 2 and 4, and fully-connected architectures, respectively.
As can be learned from Table I, our proposed algorithm offers scalability, since its computational complexity grows with lower exponents on key system parameters such as the number of users $K$, the number of RIS elements $R$ and the number of BS antennas $N$. This advantage is in addition to the superior communications performance, which shall be throughly demonstrated in the next section. For systems of smaller sizes, however, the performance advantage of the proposed method may come at the cost of a higher complexity compared to the SotA technique [26], due to the large number of iterations required for convergence of the algorithm, as shown in Figure 2. The improved sum-rate and increased robustness observed across multiple BD-RIS architectures, resulting from the algorithm’s design, are thoroughly presented and analyzed in the subsequent section.
IV Simulation Results
This section aims to showcase the effectiveness of the proposed BD-RIS-aided system in enhancing communication performance via computer simulations executed with the following parameters. I will be assumed throughout the experiments that the BS is equipped with $N=5$ TX antennas, serving $K=5$ single RX antenna users, the RIS structures consist of $R=32$ REs and the maximum transmit power $P_{\mathrm{max}}$ ranges from 0 to $20\,\mathrm{dBm}$. Furthermore, the parameters utilized in Algorithm 1 are the same as those previously used to generate the results shown in Figure 2. In addition to the performance analysis of the scattering matrix design presented in this article, the results also showcase the sum-rate behavior of our proposed scattering matrix in combination with different SotA beamforming methods [42].
In accordance with [26], the channel model adopted in this article assumes i.i.d. Rayleigh fading with distance-dependent large-scale fading expressed as $\Upsilon(d)=C_{0}(\frac{d}{d_{0}})^{-\rho}$, where $C_{0}=-30\mathrm{dB}$ is the reference channel gain at a reference distance $d_{0}=1$m, $\rho=2.2$ is the path-loss exponent, and d=50d=50m is the distance between the BS and BD-RIS  [^6] Although i.i.d fading channels are, strictly speaking, inconsistent with two-dimensional (2D) surfaces, as shown in [43], we emphasize that this particular channel model is adopted here merely in order to facilitate a fair comparison with the SotA method [26], since the scattering matrix design itself does not require any specific model to be feasible.. The system is assumed to operate at 2.42.4GHz, and each user experiences identical noise with a noise power $N_{0}=-80\mathrm{dBm}$.
The first set of results is presented in Figure 3 to demonstrate the effectiveness of the proposed method in enhancing system performance. Different architectures, such as the single-, group-, and fully-connected are considered, and results comparing the sum-rate performance achieved by the proposed and the SotA scattering matrix design [26] are provided. Notably, the results indicate that the proposed scattering matrix outperforms the SotA under the same beamforming conditions, $i.e.$, using uniform power allocation. In particular, as seen in Figure 3a, the proposed single-connected scattering matrix design surpasses both the SotA single- and group-connected $(R_{G}=2)$ architectures.
(a) Single-connected
(b) Group-connected
(c) Fully-connected
Figure 3: Comparison of sum-rate performance of the proposed vs. SotA [26] scattering matrix design with uniform power allocation, considering the single-connected “SC”, group-connected “GC”, with group sizes of 2 and 4, “GC(2)” and “GC(4)”, and the fully-connected “FC” architecture.
Figure 4: CDF of sum-rate performance of the proposed vs. SotA [26] scattering matrix design with uniform power allocation, considering the fully-connected “FC”, group-connected “GC”, with group sizes of 2 and 4, “GC(2)” and “GC(4)”, and the single-connected ”SC” architecture.
Complementing the assessment in terms of sum-rate performance, Figure 4 further shows the cumulative distribution function (CDF) of the sum-rate performance of the proposed BD-RIS scattering matrix design, alongside the SotA method [26] for all considered architectures, $i.e.$, single-, group- with group sizes of 2 and 4, and fully-connected.
The results show that, for architectures of comparable complexity, the proposed scattering matrix design consistently outperforms the SotA design. Additionally, it is observed that the single-connected architecture using the proposed method achieves performance comparable to the group-connected architecture with a groups size of 2, which has an order of magnitude higher complexity in this scenario.
Next, Figure 5 presents a comparison of the sum-rate performance of the proposed scattering matrix with different beamforming schemes, namely uniform power allocation “UP”, FP-based beamforming “BF” and power allocation “PA” [42], against the SotA scattering matrix with “UP” [26]. It is observed that the use of the FP-based TX beamformer further improves the sum-rate performance. In particular, the sum-rate achieved by the proposed single-connected architecture in combination with the SotA FP-based beamforming “BF” matrix [42] is emphasized.
Finally, Figure 6 presents the sum-rate performance of the proposed method in comparison to the SotA scattering matrix design technique of [26], in combination with uniform power allocation “UP”, as a function of the number of REs $R$. This result further builds upon the claim that for all BD-RIS architectures considered, $i.e.$, single-, group-, and fully-connected, and under every evaluated scenario the proposed scattering matrix consistently outperforms the SotA benchmark [26]. Importantly, the performance gain is maintained as the number of reflective elements increases, which highlights the robustness of the proposed approach, and showcases that the proposed solution is scalable.
(a) Single-connected
(b) Group-connected
(c) Fully-connected
Figure 5: Comparison of sum-rate performance using different beamforming schemes, considering the proposed and SotA [26] single-connected “SC”, group-connected “GC”, with group sizes of 2 and 4, “GC(2)” and “GC(4)”, and the fully-connected “FC” architecture.
Figure 6: Comparison of sum-rate performance of the proposed vs. SotA [26] scattering matrix designs as a function of the number of REs, i.e. $R$, considering the fully-connected “FC”, group-connected “GC”, with group sizes of 2 and 4, “GC(2)” and “GC(4)”, and the single-connected ”SC” architecture.
V Conclusion
A novel sum-rate maximization scheme for RBD-RIS-aided MU-MISO systems is presented. To this end, an optimization problem was formulated to determine the BD-RIS scattering matrix configuration that maximizes the users’ sum-rate, and solved using a Riemannian manifold steepest descent method with a closed-form gradient. Simulation results validated the effectiveness of the proposed approach, showing consistent performance gains over the SotA, where the improvements arise from jointly exploiting the geometric structure of the manifold and the direct sum-rate based problem formulation. Overall, the results show that manifold-based scattering matrix optimization enables scalable and robust performance in RIS-aided systems. Future work will extend the framework to include transmit beamforming design.
Appendix A Proof of Lemma 1
Lemma 1.
Consider $\mathbf{A}\in\mathbb{C}^{N\times N}$ with entries defined as
$\mathbf{A}=\begin{pmatrix}a_{1,1}&a_{1,2}&\cdots&a_{1,N}\\ a_{2,1}&a_{2,2}&\cdots&a_{2,N}\\ \vdots&\vdots&\ddots&\vdots\\ a_{N,1}&a_{N,2}&\cdots&a_{N,N}\end{pmatrix}$. (46)
Then
$\nabla_{\mathbf{A}}\mathrm{Tr}\{\mathbf{A}^{\star}\mathbf{A}\}=2\frac{\partial\mathrm{Tr}\{\mathbf{A}^{\star}\mathbf{A}\}}{\partial\mathbf{A}^{\star}}=2\mathbf{A}^{T},$ (47)
where
$\mathrm{Tr}\{\mathbf{A}^{\star}\mathbf{A}\}=\sum_{i=1}^{N}\sum_{k=1}^{N}a_{i,k}^{\star}a_{k,i},$ (48)
with $a_{i,k}^{\star}$ and $a_{k,i}\in\mathbb{C}$, such that they can be written in the form $x=\Re{x}+j\Im{x}$.
Proof.
For simplicity, let $\phi=\mathrm{Tr}\{\mathbf{A}^{\star}\mathbf{A}\}$. Then,
$\frac{\partial\phi}{\partial\Re{\mathbf{A}}}$ $=\begin{pmatrix}\frac{\partial\phi}{\partial\Re a_{1,1}}&\frac{\partial\phi}{\partial\Re a_{1,2}}&\cdots&\frac{\partial\phi}{\partial\Re a_{1,N}}\\ \frac{\partial\phi}{\partial\Re a_{2,1}}&\frac{\partial\phi}{\partial\Re a_{2,2}}&\cdots&\frac{\partial\phi}{\partial\Re a_{2,N}}\\ \vdots&\vdots&\ddots&\vdots\\ \frac{\partial\phi}{\partial\Re a_{N,1}}&\frac{\partial\phi}{\partial\Re a_{N,2}}&\cdots&\frac{\partial\phi}{\partial\Re a_{N,N}}\end{pmatrix}$ (49)
$=2\Re{\begin{pmatrix}a_{1,1}&a_{2,1}&\cdots&a_{N,1}\\ a_{1,2}&a_{2,2}&\cdots&a_{N,2}\\ \vdots&\vdots&\ddots&\vdots\\ a_{1,N}&a_{2,N}&\cdots&a_{N,N}\end{pmatrix}}$
and
$j\frac{\partial\phi}{\partial\Im{\mathbf{A}}}$ $=\begin{pmatrix}\frac{\partial\phi}{\partial\Im a_{1,1}}&\frac{\partial\phi}{\partial\Im a_{1,2}}&\cdots&\frac{\partial\phi}{\partial\Im a_{1,N}}\\ \frac{\partial\phi}{\partial\Im a_{2,1}}&\frac{\partial\phi}{\partial\Im a_{2,2}}&\cdots&\frac{\partial\phi}{\partial\Im a_{2,N}}\\ \vdots&\vdots&\ddots&\vdots\\ \frac{\partial\phi}{\partial\Im a_{N,1}}&\frac{\partial\phi}{\partial\Im a_{N,2}}&\cdots&\frac{\partial\phi}{\partial\Im a_{N,N}}\end{pmatrix}$ (50)
$=2j\Im{\begin{pmatrix}a_{1,1}&a_{2,1}&\cdots&a_{N,1}\\ a_{1,2}&a_{2,2}&\cdots&a_{N,2}\\ \vdots&\vdots&\ddots&\vdots\\ a_{1,N}&a_{2,N}&\cdots&a_{N,N}\end{pmatrix}}$,
where it is clear that
$\frac{\partial\phi}{\partial\Re\mathbf{A}}=2\Re\mathbf{A}^{T},$ (51)
and
$j\frac{\partial\phi}{\partial\Im\mathbf{A}}=2j\Im\mathbf{A}^{T}.$ (52)
Straightforwardly, the Generalized Complex Derivative and Conjugate Complex Derivative [38] are obtained as
$\frac{\mathrm{d}\phi}{\mathrm{d}\mathbf{A}}=\frac{1}{2}\Big(2\Re\mathbf{A}^{T}-2j\Im\mathbf{A}^{T}\Big)=\mathbf{A}^{H},$ (53)
$\frac{\mathrm{d}\phi}{\mathrm{d}\mathbf{A}^{\star}}=\frac{1}{2}\Big(2\Re\mathbf{A}^{T}+2j\Im\mathbf{A}^{T}\Big)=\mathbf{A}^{T}.$ (54)
As evidence that $\phi$ is a real-valued function, the following is provided
$\mathrm{Tr}\{\mathbf{A}^{\star}\mathbf{A}\}=\sum_{i}^{N}|a_{i,i}|^{2}+2\Re\{\sum_{i}^{N}\sum_{k=i+1}^{N}a_{i,k}^{\star}a_{k,i}\},$ (55)
as such, the Complex Gradient Matrix [38] can be computed as follows
$\nabla_{\mathbf{A}}\mathrm{Tr}\{\mathbf{A}^{\star}\mathbf{A}\}=2\frac{\partial\mathrm{Tr}\{\mathbf{A}^{\star}\mathbf{A}\}}{\partial\mathbf{A}^{\star}}=2\mathbf{A}^{T},$ (56)
hence, (24) is proven.
Appendix B Proof of Lemma 2
Lemma 2.
Consider $\mathbf{a}\in\mathbb{C}^{K\times 1}$, $\mathbf{B}\in\mathbb{C}^{K\times K}$, and $\mathbf{c}\in\mathbb{C}^{K\times 1}$, such that $d=\mathbf{a}^{T}\mathbf{B}\mathbf{c}$, where $d\in\mathbb{C}$ and
$d=\left(\!\begin{array}[]{c}a_{1}\\ a_{2}\\ \vdots\\ a_{N}\end{array}\!\right)^{T}\!\!\!\left(\!\begin{array}[]{cccc}B_{1,1}&B_{1,2}&\cdots&B_{1,N}\\ B_{2,1}&B_{2,2}&\cdots&B_{2,N}\\ \vdots&\vdots&\ddots&\vdots\\ B_{N,1}&B_{N,2}&\cdots&B_{N,N}\end{array}\!\right)\!\!\!\left(\!\begin{array}[]{c}c_{1}\\ c_{2}\\ \vdots\\ c_{N}\end{array}\!\right).$ (57)
Consider the function $f(d)=|d|^{2}$. The gradient of $f(d)$ with respect to $\mathbf{B}$ is given as
$\nabla_{\mathbf{B}}|d|^{2}=d\mathbf{a}^{\star}\mathbf{c}^{H},$ (58)
where $\mathbf{a}=\Re\mathbf{a}+j\Im\mathbf{a}$, $\mathbf{B}=\Re\mathbf{B}+j\Im\mathbf{B}$ and $\mathbf{c}=\Re\mathbf{c}+j\Im\mathbf{c}$, where $\Re\backslash\Im\mathbf{a}$, $\Re\backslash\Im\mathbf{c}\in\mathbb{R}^{K\times 1}$ and $\Re\backslash\Im\mathbf{B}\in\mathbb{R}^{K\times K}$.
Proof.
Following the steps for calculating the gradient of complex valued functions, as described in [38]. The derivatives with respect to the real and imaginary parts of $\mathbf{B}$ are calculated, as shown below in
$\frac{\partial|d|^{2}}{\partial\Re\mathbf{B}}=\frac{\partial|d|^{2}}{\partial d}\frac{\partial d}{\partial\Re\mathbf{B}}+\frac{\partial|d|^{2}}{\partial d^{\star}}\frac{\partial d^{\star}}{\partial\Re\mathbf{B}},$ (59)
and
$\frac{\partial|d|^{2}}{\partial\Im\mathbf{B}}=\frac{\partial|d|^{2}}{\partial d}\frac{\partial d}{\partial\Im\mathbf{B}}+\frac{\partial|d|^{2}}{\partial d^{\star}}\frac{\partial d^{\star}}{\partial\Im\mathbf{B}}.$ (60)
Straightforwardly
$\frac{\partial|d|^{2}}{\partial d}=d^{\star},$ (61)
$\frac{\partial|d|^{2}}{\partial d^{\star}}=d.$ (62)
Expanding $d=\mathbf{a}^{T}\mathbf{B}\mathbf{c}$ and $d^{\star}=\mathbf{c}^{H}\mathbf{B}^{H}\mathbf{a}^{\star}$ as a product of complex numbers. The following derivatives are obtained
$\frac{\partial d}{\partial\Re\mathbf{B}}$ $=\begin{pmatrix}\frac{\partial d}{\partial\Re B_{1,1}}&\frac{\partial d}{\partial\Re B_{1,2}}&\cdots&\frac{\partial d}{\partial\Re B_{1,N}}\\ \frac{\partial d}{\partial\Re B_{2,1}}&\frac{\partial d}{\partial\Re B_{2,2}}&\cdots&\frac{\partial d}{\partial\Re B_{2,N}}\\ \vdots&\vdots&\ddots&\vdots\\ \frac{\partial d}{\partial\Re B_{N,1}}&\frac{\partial d}{\partial\Re B_{N,2}}&\cdots&\frac{\partial d}{\partial\Re B_{N,N}}\end{pmatrix}$ (63)
$={\begin{pmatrix}a_{1}c_{1}&a_{1}c_{2}&\cdots&a_{1}c_{N}\\ a_{2}c_{1}&a_{2}c_{2}&\cdots&a_{2}c_{N}\\ \vdots&\vdots&\ddots&\vdots\\ a_{N}c_{1}&a_{N}c_{2}&\cdots&a_{N}c_{N}\end{pmatrix}}=\mathbf{a}\mathbf{c}^{T},$
$\frac{\partial d}{\partial\Im\mathbf{B}}$ $=\begin{pmatrix}\frac{\partial d}{\partial\Im B_{1,1}}&\frac{\partial d}{\partial\Im B_{1,2}}&\cdots&\frac{\partial d}{\partial\Im B_{1,N}}\\ \frac{\partial d}{\partial\Im B_{2,1}}&\frac{\partial d}{\partial\Im B_{2,2}}&\cdots&\frac{\partial d}{\partial\Im B_{2,N}}\\ \vdots&\vdots&\ddots&\vdots\\ \frac{\partial d}{\partial\Im B_{N,1}}&\frac{\partial d}{\partial\Im B_{N,2}}&\cdots&\frac{\partial d}{\partial\Im B_{N,N}}\end{pmatrix}$ (64)
$=j{\begin{pmatrix}a_{1}c_{1}&a_{1}c_{2}&\cdots&a_{1}c_{N}\\ a_{2}c_{1}&a_{2}c_{2}&\cdots&a_{2}c_{N}\\ \vdots&\vdots&\ddots&\vdots\\ a_{N}c_{1}&a_{N}c_{2}&\cdots&a_{N}c_{N}\end{pmatrix}}=j\mathbf{a}\mathbf{c}^{T}.$
Similarly, for $d^{\star}$
$\frac{\partial d^{\star}}{\partial\Re\mathbf{B}}=\mathbf{a}^{\star}\mathbf{c}^{H},$ (65)
$\frac{\partial d^{\star}}{\partial\Im\mathbf{B}}=-j\mathbf{a}^{\star}\mathbf{c}^{H}.$ (66)
Finally
$\frac{\partial|d|^{2}}{\partial\Re\mathbf{B}}=d^{\star}\mathbf{a}\mathbf{c}^{T}+d\mathbf{a}^{\star}\mathbf{c}^{H},$ (67)
$\frac{\partial|d|^{2}}{\partial\Im\mathbf{B}}=jd^{\star}\mathbf{a}\mathbf{c}^{T}-jd\mathbf{a}^{\star}\mathbf{c}^{H}.$ (68)
$\frac{\mathrm{d}|d|^{2}}{\mathrm{d}\mathbf{B}^{\star}}=\frac{1}{2}\Bigg(\frac{\partial|d|^{2}}{\partial\Re\mathbf{B}}+j\frac{\partial|d|^{2}}{\partial\Im\mathbf{B}}\Bigg)=d\mathbf{a}^{\star}\mathbf{c}^{H}.$ (69)
As the absolute value squared of a complex number is a real value, $i.e.$, $|d|^{2}$ is a real function, the gradient of $f(d)$ with respect to $\mathbf{B}$ is given as
$\nabla_{\mathbf{B}}|d|^{2}=2\frac{\mathrm{d}|d|^{2}}{\mathrm{d}\mathbf{B}^{\star}}=2d\mathbf{a}^{\star}\mathbf{c}^{H}=2\big(d^{\star}\mathbf{a}\mathbf{c}^{T}\big)^{\star},$ (70)
hence, (28) and (29) are proven. ∎