### v4

# Reciprocal Beyond-Diagonal Reconfigurable Intelligent Surface  (BD-RIS):  Scattering Matrix Design via Manifold Optimization

Abstract
Beyond-diagonal reconfigurable intelligent surfaces (BD-RISs) are emerging as a transformative technology in wireless communications, enabling enhanced performance and quality of service (QoS) of wireless systems in harsh urban environments due to their relatively low cost and advanced signal processing capabilities. Generally, BD-RIS systems are employed to improve robustness, increase achievable rates, and enhance energy efficiency of wireless systems in both direct and indirect ways. The direct way is to produce a favorable propagation environment via the design of optimized scattering matrices, while the indirect way is to reap additional improvements via the design of multiple-input multiple-output (MIMO) beamformers that further exploit the latter “engineered” medium. In this article, the problem of sum-rate maximization via BD-RIS is examined, with a focus on feasibility, namely low-complexity physical implementation, by enforcing reciprocity in the BD-RIS design in a manner that adheres to the geometry of the manifold of symmetric matrices. To that end, the sum-rate objective is transformed into a quadratic function via fractional programming (FP), augmented via the also quadratic reciprocity constraint in the form of a regularization term, while the unitary constraint is dealt with via a manifold optimization framework. Simulation results demonstrate the effectiveness of the proposed method in outperforming current state-of-the-art (SotA) approaches in terms of sum-rate maximization.
I Introduction
Intelligent metasurfaces, or reconfigurable intelligent surfaces (RISs), are an emerging resource in the area of wireless communications, offering a new paradigm for controlling foundational features of electromagnetic fields to enhance system performance [2, 3, 4, 5, 6, 7]. These surfaces consist of individual elements, $i.e.$, passive circuits, with controllable properties that allow them to filter and reradiate the incoming signals without amplification, which is accomplished by implementing delays, reducing the amplitude, or changing wave polarization via the circuit’s local reprogrammable impedance.
Thanks to their low-power and reconfigurable nature, the so-called wave-domain signal processing approach offered by RISs offer several key advantages over classical analog/digital signal processing, including ease of deployment and substantial reductions in end-to-end latency and processing delay [8].
Motivated by these advantages, RISs have attracted enormous attention as an effective solution for enhancing wireless channel quality and consequently the communication performance of multiple-input single-output (MISO), multiple-input multiple-output (MIMO), and integrated sensing and communications (ISAC) systems [2]. Indeed, intelligent surfaces revolutionize wireless systems design by allowing dynamic reconfiguring of the propagation environment, achieving substantial gains without the need for additional active radio chains [9, 10].
Early research on the approach predominantly focused on diagonal reconfigurable intelligent surface (D-RIS) architectures, whose scattering matrices enabled simpler analysis and straightforward implementation [11, 12, 13, 14, 15, 16, 17]. Their use has been extensively studied across various communication settings under various design objectives, including sum-rate maximization [11], spectral efficiency improvement [12], energy efficiency enhancement [13], transmit power reduction [14], and physical layer security [15]. To mention a few illustrative examples, in [16] a weighted sum-rate maximization problem under perfect and imperfect channel state information (CSI) was proposed, with algorithms developed for each case. Alternatively, in [17], an active RIS-aided downlink rate-splitting multiple access system is proposed, and its sum-rate performance was analyzed via a novel algorithm. To this extent, a joint optimization problem was formulated to design the transmit (TX) beamforming and rate allocation vectors together with the RIS scattering matrix. The approach involves a two-stage design, where fractional programming (FP) techniques were used in combination with quadratic constraint quadratic programming (QCQP) methods to optimize the respective variables. Another excellent example is [15], whereby the authors utilize the D-RIS in an ISAC framework, applying manifold optimization to jointly improve sum-rate performance and enhance security against eavesdropping.
Although the conventional diagonal architecture offers analytical convenience, it also imposes structural constraints that limit the degrees-of-freedom (DoF), restricting its ability to manipulate the signal space. To overcome these limitations, subsequent studies have focused on beyond-diagonal reconfigurable intelligent surface (BD-RIS) architectures, which significantly increase the DoF available for signal shaping [18, 19, 20, 21, 22, 23, 24, 25, 26]. However, the increase in DoF, achieved through enhanced connectivity of the BD-RIS, is generally accompanied by higher hardware complexity and implementation cost. As a result, a substantial body of literature has investigated the trade-off between achievable performance and hardware complexity in BD-RIS architectures [27, 28].
In this context, advanced optimization techniques have been employed to effectively exploit the additional DoF while respecting structural constraints. Both FP and first-order methods, including projected gradient and accelerated gradient schemes [29], have been widely adopted in RIS optimization problems to improve convergence speed and computational efficiency. While these approaches typically enforce constraints through Euclidean projections onto feasible sets, manifold optimization instead leverages the intrinsic geometry of the constraint space. By operating directly on smooth manifolds via tangent-space projections and retraction steps [30, 31], it maintains feasibility along the optimization trajectory while explicitly accounting for the geometry of the constraint set. Junior et al. [18] provide a comprehensive tutorial on manifold optimization techniques and their application in RIS-aided massive MIMO (mMIMO) systems, where it is shown that adhering to the manifold geometry of the constraints consistently yields better solutions than standard optimization methods, which motivates the approach adopted in this article. Several recent works have applied these principles to specific BD-RIS design problems, employing manifold optimization to enhance secrecy rate [19] and to jointly configure the scattering matrix and TX precoder [20]. Both these as well as similar studies such as [32, 33], however, are conducted for non-reciprocal settings, without enforcing symmetry.
Enforcing a symmetric scattering matrix ensures compatibility with physically realizable passive architectures, avoiding the need for active components, magnetic biasing, or non-reciprocal circuit elements, which increase hardware complexity, power consumption, and implementation cost [7, 24]. One way to enforce the unitary and symmetry constraints inherent to reciprocal BD-RIS is through a penalty dual decomposition (PDD) framework, as proposed in [23], where the scattering matrix and transmit beamformers are jointly optimized for both sum-rate maximization and transmit power minimization problems. While effective, the approach relies on auxiliary-variable decoupling and a double-loop procedure, is primarily tailored to fully- and group-connected BD-RIS architectures, and incurs relatively high computational complexity. To address these limitations, [34] develops an architecture-independent framework based on partially proximal alternating direction method of multipliers (pp-ADMM), which reformulates the BD-RIS constraints via the admittance matrix and enables efficient closed-form or low-complexity updates at each iteration. This general and computationally efficient framework is adopted in this work as a benchmark for performance and complexity comparisons.
Alternative BD-RIS design strategies extend beyond direct optimization from the sum-rate objective and instead focus on hybrid approaches. For example, in [21], two methods for designing BD-RIS scattering matrices are proposed. The first method seeks to maximize the total signal power received by the users following a passive maximum ratio transmission (MRT) strategy, while the second approach aims at nullifying the multi-user interference via a passive zero forcing (ZF) strategy. In contrast, this work adopts a fundamentally different strategy: rather than decoupling the design into separate MRT or ZF criteria, the scattering matrix is optimized with respect to the sum-rate objective. This direct formulation accounts for both signal enhancement and interference suppression, enabling these aspects to be addressed simultaneously within a unified and coherent optimization framework.
Another notable example is presented by Fang et al. [22], where the authors propose an alternative methodology—leveraging convex relaxation techniques—to address the non-convex BD-RIS problem. Specifically, they reformulate the non-convex constraints into a convex set, enabling efficient solution of the relaxed problem. The solution is then mapped back to the feasible region of the original non-convex set using their proposed symmetric unitary projection.
More recently, research has shifted toward non-reciprocal BD-RIS (NRBD-RIS) designs, particularly within a full-duplex framework [24, 25]. For example, Liu et al. [24] investigate their application in wireless circulator systems, demonstrating that properly optimized non-reciprocal scattering matrices yield improvements in both spectral efficiency and interference management compared to existing reciprocal structures, further expanding the applicability of BD-RIS in wireless networks. However, the practical realization of non-reciprocal BD-RIS entails increased hardware complexity and resource requirements, which may limit their suitability in large-scale or cost-sensitive deployments. Although non-reciprocal designs can be advantageous in specific scenarios, reciprocal BD-RIS architectures remain highly relevant for the broader range of applications where simplicity, scalability, and energy efficiency are essential.
All in all, we recognize the maximization of sum-rate via the design of reciprocal BD-RIS scattering matrices as a dual manifold problem which, to the best of our knowledge, has not yet been addressed as such via an adequate manifold optimization method, likely due to the challenge of dealing with two manifolds of different geometries simultaneously. Recently, this challenge was addressed in [35], where a Riemannian optimization algorithm operating on the solution space defined by the intersection manifold of unitary and symmetric matrices was proposed. [^1] We emphasize that a pre-print of [35] first appeared only in Jan. 2026 https://arxiv.org/abs/2601.13877, after the submission of this article which was available at http://www.arxiv.org/abs/2509.20246 since Sep. 2025. That formulation is, however, restricted to the fully-connected architecture and primarily considers the single-user case, where inter-user interference is absent.
In contrast to the latter, in this article, the dual-manifold nature of the reciprocal BD-RIS scattering matrix design problem is addressed under a general setting [^2] For the sake of simplicity and to allow direct comparison with state-of-the-art (SotA) methods we explicitly focus on multi-user (MU)-MISO systems, for which closed-form gradients are derived, while emphasizing that the proposed method applies directly to the general MU-MIMO case. See Subsection II-C for more details. by incorporating the reciprocity constraint as a quadratic regulation term into the sum-rate objective transformed also into a quadratic function via FP, while enforcing the unitarity constraint by solving the resulting regularized problem via manifold optimization over the Stiefel manifold. Our main contributions are summarized as follows:
  • Novel technique with improved performance: A novel method is proposed to design the scattering matrix of reciprocal BD-RIS (RBD-RIS), under which both the symmetry and unitary constraints are sensibly enforced, in a manner that accommodates both manifold geometries. In particular, in the proposed method, the quadratic geometry of the symmetry constraint manifold is incorporated into the quadratic sum-rate objective reformulated via FP, while the geometry of the Stiefel unitary constraint manifold is dealt with via manifold optimization. The new method, which has complexity similar to that of SotA techniques, is shown via simulations to outperform the latter under equivalent conditions.
  • Original closed-form expressions: While the method is applicable to general MU-MIMO systems, a closed-form expression for the gradient of the proposed regularized and FP-reformulated objective function is provided for the MU-MISO setup, enabling the implementation of the method via conjugate gradient ascent (CGA) in that case.
  • High hardware compatibility: Thanks to the above, the proposed scattering matrix design of RBD-RIS under MU-MISO settings can be implemented in a highly controlled manner that facilitates efficient hardware design [36]. In particular, with the exception of an Armijo line search, the key steps of the method, namely, the tangential projection, gradient evaluation, and updating of FP variables, are achieved via successive evaluation of closed-form expressions, enabling dedicated signal processing hardware to be optimized. [^3] It is a well-known and widely recognized principle in digital signal processing (DSP) and very large scale integration (VLSI) designs that closed-form algorithms reduce to a fixed sequence of arithmetic operations, favoring the stability, reliability, power consumption and efficacy of hardware implementation, as a result of the corresponding deterministic pipelines, parallelization and static scheduling [36].
Organization: Section II introduces the BD-RIS-aided MU-MISO system model, presents the different architectures considered, and formulates the problem addressed in the article. Next, a general solution to the first stage of the optimization problem, namely the scattering matrix design, is provided in Section III. Simulation results are offered in Section IV, which help evaluate the performance of the proposed designs. Finally, a brief conclusion is given in Section V, followed by Appendix A and B, where proofs of two Lemmas are detailed.
Notation: Unless otherwise specified, $\mathbf{X}$ and $\mathbf{x}$ denote matrices and vectors. The absolute value, $\ell^{2}$ and Frobenius norm are denoted by $|\cdot|$, $\|\cdot\|_{2}$, $\|\cdot\|_{F}$. The transpose and the Hermitian transpose are denoted by $(\cdot)^{T}$ and $(\cdot)^{H}$. In addition, $\mathbf{X}_{i,j}$ denotes the $i$-th row and $j$-th column element of the matrix $\mathbf{X}$, $\mathbf{X}_{i:\bar{i},j:\bar{j}}$ extracts the elements from the $i$-th to the $\bar{i}$-th row, and $j$-th to $\bar{j}$ column from the matrix $\mathbf{X}$. Furthermore, $\text{diag}(\mathbf{x})=(x_{1},x_{2},\dots,x_{k})$ denotes a diagonal matrix, where the main diagonal is $\mathbf{x}$, and $\text{blkdiag}\left(\mathbf{X}_{1},\mathbf{X}_{2},\dots,\mathbf{X}_{k}\right)$ denotes a $G\times G$ block diagonal matrix with off diagonal elements 0, where the blocks are $\mathbf{X}_{1},\mathbf{X}_{2},\dots,\mathbf{X}_{k}$. The sets of complex and real numbers are denoted by $\mathbb{C}$ and $\mathbb{R}$, while the real and imaginary values of a complex number are denoted by $\Re\{\cdot\}$ and $\Im\{\cdot\}$, respectively. The circular-symmetric complex normal random variable with $\sigma^{2}$ variance is denoted as $\mathcal{CN}(0,\sigma^{2})$.
II System Model
Consider a BD-RIS aided downlink MU-MISO system as illustrated in Figure 1, where a base station (BS) with $N$ TX antennas serves $K$ single receive (RX) antenna users [^4] This scenario is considered due to its analytical tractability and simplicity, while enabling direct comparisons with SotA alternatives. The methods can, however, be easily generalized to MU-MIMO settings, as described in Subsection II-C. with the help of a BD-RIS consisting of $R$ reflective elements (REs), The associated BD-RIS scattering matrix is represented by $\mathbf{\Theta}\in\mathbb{C}^{R\times R}$. The channel linking the BS and the BD-RIS ($i.e.$, the BS-to-BD-RIS channel) is denoted as $\mathbf{H}_{\mathrm{TX}}\in\mathbb{C}^{R\times N}$, and the channel vector connecting the BD-RIS and the $k$-th user is $\mathbf{h}_{k}\in\mathbb{C}^{R\times 1}$.
The transmit signal vector is given by $\mathbf{x}=\mathbf{V}\mathbf{s}$, where the information symbols $\mathbf{s}\in\mathbb{C}^{K\times 1}$ satisfy $\mathbb{E}[\mathbf{s}\mathbf{s}^{H}]=\mathbf{I}$, and the beamforming matrix $\mathbf{V}\in\mathbb{C}^{N\times K}$ meets the power constraint $\|\mathbf{V}\|_{F}^{2}\leq P_{\mathrm{max}}$, with $P_{\mathrm{max}}$ being the maximum transmit power at the BS. It is assumed that the channel between the BS and the $k$-th user is blocked [^5] This assumption captures the most challenging and conceptually relevant scenario, where the impact and benefits of RIS-assisted propagation are most clearly observed. In particular, it highlights coverage extension and blockage mitigation, while reducing the need for network-level optimization measures such as deploying additional base stations. As a result, the blocked-user assumption is widely adopted in RIS-related studies, enabling direct comparison of the proposed technique with SotA methods, $e.g.$ [21, 34]. Importantly, this assumption is not required by our proposed approach, since the scattering matrix design depends only on the cascaded BS-to-RIS-to-user channels. Nevertheless, the overall performance under this setting relies on the BS-to-RIS channel having a sufficiently high rank, which holds only in rich-scattering, large-antenna setups, motivating an extension of the work under poorly scattered channels with line-of-sight (LoS) links between the BS and users. Such an extension also requires, however, further work on the beamforming algorithm, thus falling beyond the scope of this article, and therefore will be pursued in a follow-up contribution., such that a focus is placed on the performance gains due to the BD-RIS. Thus, the signal $r_{k}\in\mathbb{C}$ received at user $k$ can be expressed as
$r_{k}=\mathbf{h}_{k}^{T}\mathbf{\Theta}\mathbf{H}_{\mathrm{TX}}\mathbf{x}+n_{k},$ (1)
where $n_{k}\sim\mathcal{CN}(0,N_{0})$ denotes additive white Gaussian noise (AWGN) with power $N_{0}$.
Figure 1: Illustration of the system model, where a BS with $N$ TX antennas serves $K$ single-antenna users through an $R$-Element RBD-RIS, without a LoS link2 between the BS and the users.
The compact vector representation in (1) is inspired by the formulation used in [21], and is given as
$\mathbf{r}=[r_{1},r_{2},\dots,r_{K}]^{T},$ (2a)
$\mathbf{n}=[n_{1},n_{2},\dots,n_{K}]^{T},$ (2b)
$\mathbf{V}=[\mathbf{v}_{1},\mathbf{v}_{2},\dots,\mathbf{v}_{K}],$ (2c)
$\mathbf{H}_{\mathrm{RX}}=[\mathbf{h}_{1},\mathbf{h}_{2},\dots,\mathbf{h}_{K}]^{T},$ (2d)
$\mathbf{H}_{\mathrm{TX}}=\left[\mathbf{w}_{1},\mathbf{w}_{2},\dots,\mathbf{w}_{N}\right],$ (2e)
$\mathbf{r}=\mathbf{H}_{\mathrm{RX}}\boldsymbol{\Theta}\mathbf{H}_{\mathrm{TX}}\mathbf{x}+\mathbf{n},$ (2f)
where $\mathbf{r}\in\mathbb{C}^{K\times 1}$, $\mathbf{H}_{\mathrm{RX}}\in\mathbb{C}^{K\times R}$, $\mathbf{w}_{n}\in\mathbb{C}^{R\times 1}$ with $n\in\{1,2,\dots,N\}$, and $\mathbf{v}_{m}\in\mathbb{C}^{N\times 1}$ with $m\in\{1,2,\dots,K\}$, denote the received signal vector, BD-RIS-to-user $k$ channel matrix, BS-to-BD-RIS channel vector, and beamforming vector, respectively.
II-A Scattering Matrix Definition
Depending on the reconfigurable impedance network of the BD-RIS, three different architectures are considered, namely:
  1. 1.
    Single-connected BD-RIS: Equivalent to the conventional D-RIS, where each RE is independently connected to a single grounded reconfigurable impedance with no interconnections among REs. Consequently, the resulting scattering matrix is inherently diagonal, which naturally satisfies the symmetry constraints, while also being unitary due to the lossless nature [^6] According to microwave network theory, a lossless reconfigurable impedance network must satisfy a unitary scattering matrix condition, thereby ensuring energy conservation [4]. Furthermore, the reciprocity property, which enforces symmetry of the BD-RIS scattering matrix, follows directly from electromagnetic reciprocity and microwave network theory. In particular, any passive, linear and time-invariant structure exhibits symmetric transmission characteristics, $i.e.$, the coupling between any two ports of the BD-RIS is identical in both directions, which is mathematically captured by the symmetric scattering matrix [37].. The aforementioned constraints are described as
    $\mathcal{S}_{\text{SC}_{1}}=\left\{\mathbf{\Theta}:[\mathbf{\Theta}]_{i,j}=0,\ \forall i\neq j\right\},$ (3)
    $\mathcal{S}_{\text{SC}_{2}}=\left\{\mathbf{\Theta}:\left|[\mathbf{\Theta}]_{i,j}\right|=1,\ \forall i=j\right\},$ (4)
    where $i,j\in\{1,2,\dots,R\}$.
  2. 2.
    Fully-connected BD-RIS: In this architecture, every RE is connected to all others via a reconfigurable impedance network, allowing a wave incident on one element to be reflected by all others. As a result, the corresponding scattering matrix is full and—similarly to the other architectures—must satisfy both the symmetry and the unitary constraint, which for this case are given as
    $\mathcal{S}_{\text{FC}_{1}}=\left\{\mathbf{\Theta}:\mathbf{\Theta}=\mathbf{\Theta}^{T}\right\},$ (5)
    $\mathcal{S}_{\text{FC}_{2}}=\left\{\mathbf{\Theta}:\mathbf{\Theta}\mathbf{\Theta}^{H}=\mathbf{I}\right\}.$ (6)
  3. 3.
    Group-connected BD-RIS: In this intermediate architecture the REs are partitioned into $G$ disjoint groups, each consisting of $R_{G}=\frac{R}{G}$ elements. Within each group, all REs are interconnected, enabling waves impinging on one element to be reflected by all others in the same group. Apart from the intra-group connections, no other link exists. Consequently, the scattering matrix can be represented as a block-diagonal matrix by ordering its elements group-by-group, such that each block corresponds to one group and must independently satisfy the unitary and symmetry constraints. Notably, the group-connected architecture generalizes the previous two cases, reducing to the fully-connected case when $G=1$ and to the single-connected case when $G=R$. Denoting the scattering matrix of groups $g\in\{1,2,\ldots,G\}$ by $\mathbf{\Theta}_{g}\in\mathbb{C}^{R_{G}\times R_{G}}$, the overall structure of the group-connected architecture is expressed by
    $\mathbf{\Theta}=\mathrm{blkdiag}(\mathbf{\Theta}_{1},\ldots,\mathbf{\Theta}_{G}),$ (7)
    and the constraints on the group-connected BD-RIS scattering matrix are given by
    $\hskip-21.52771pt{\mathcal{S}_{\text{GC}_{1}}\!\!=\!\left\{\mathbf{\Theta}:\mathbf{\Theta}_{g}\!=\!\mathbf{\Theta}_{g}^{T},\forall g\right\}\!,\!\!\!}$ (8)
    $\mathcal{S}_{\text{GC}_{2}}=\left\{\mathbf{\Theta}:\mathbf{\Theta}_{g}\mathbf{\Theta}_{g}^{H}=\mathbf{I}_{R_{G}},\;\forall g\right\}.$ (9)
    Here $\mathbf{\Theta}_{g}$ is formally defined as
    $\mathbf{\Theta}_{g}=\mathbf{\Theta}_{[R_{G}(g-1)+1\,:\,gR_{G},\;\;R_{G}(g-1)+1\,:\,gR_{G}]},$ (10)
    where $g\in\{1,2,\dots,G\}$. Figure 2, seen at the top of the next page, provides an illustrative example demonstrating the impact of beyond-diagonal connectivity on the scattering behavior of the BD-RIS.
(a) Single-Connected
(b) Group-Connected (group size 4)
(c) Fully-connected
Figure 2: Beampattern illustrating the responses of a BD-RIS with different levels of connectivity. The plots are shown in terms of the auxiliary quantities $u\triangleq\sin(\phi)\cos(\theta)$ and $v\triangleq\sin(\phi)\sin(\theta)$, where $\phi$ and $\theta$ are the azimuth and elevation angles, swept in the ranges $\pi,\pi]$ and $\pi/2,\pi/2]$, respectively. The blue lines correspond to the directions of two users, located at $(\phi_{1},\theta_{1})=(-40^{\circ},10^{\circ})$ and $(\phi_{2},\theta_{2})=(-20^{\circ},10^{\circ})$.
II-B Problem Formulation
Given $\mathbf{\Theta}$ and $\mathbf{V}$, and assuming perfect CSI  [^7] Although idealized, the assumption of perfect CSI is widely adopted in the RIS literature [14, 21, 34], as it enables evaluation of the “best-case” performance of the proposed algorithm, while imperfect CSI can only degrade performance. Moreover, in block-static channel conditions where the channel remains constant over the data transmission interval, the channel coefficients can, in principle, be estimated with arbitrarily high accuracy given sufficient training resources. In such settings, performance analysis under perfect CSI is consistent with the standard block-fading framework, where coding is performed over long coherence intervals. In contrast, in rapidly varying fading environments, channel estimation errors become unavoidable and must be explicitly incorporated into the system design. Extensions to robust formulations under imperfect CSI are feasible and have been extensively studied in the literature (see, e.g., [38, 39]). However, such developments fall beyond the scope of the present work., an achievable rate at the $k$-th user is $\log_{2}(1+\gamma_{k})$, where the signal to interference-plus-noise ratio (SINR) [^8] For the extension to the general MU-MIMO case, the SINR in (11) would need to be revisited to incorporate receive combining at the user side. is expressed as
$\gamma_{k}=\frac{\left|\mathbf{h}_{k}^{T}\mathbf{\Theta}\mathbf{H}_{\mathrm{TX}}\mathbf{v}_{k}\right|^{2}}{\sum\limits_{i\neq k}\left|\mathbf{h}_{k}^{T}\mathbf{\Theta}\mathbf{H}_{\mathrm{TX}}\mathbf{v}_{i}\right|^{2}+N_{0}}\ .\vskip-4.30554pt$ (11)
The corresponding sum-rate maximization problem is then
$(\text{P1}):\quad\underset{\mathbf{V},\mathbf{\Theta}}{\mathrm{maximize}}\quad$ $\sum_{k}\log_{2}(1+\gamma_{k})$ (12a)
subject to $\|\mathbf{V}\|_{F}^{2}\leq P_{\mathrm{max}},$ (12b)
$\mathbf{\Theta}\in\mathcal{S}_{a_{1}},$ (12c)
$\mathbf{\Theta}\in\mathcal{S}_{a_{2}},$ (12d)
where the constraint (12b) corresponds to the total transmit power limitation at the BS, ensuring that the Frobenius norm of the beamforming matrix $\mathbf{V}$ does not exceed the maximum allowable transmit power $P_{\mathrm{max}}$, $a\in\{\text{SC},\text{FC},\text{GC}\}$ denotes the connectivity level of the BD-RIS, and $\mathcal{S}_{a_{1}}$ and $\mathcal{S}_{a_{2}}$ represent the sets of scattering matrices that satisfy the reciprocal and lossless constraints, respectively.
The optimization problem (P1) is inherently non-convex due to the objective function in (12a), and the strong coupling between the TX beamforming and the scattering matrix. Additionally, the constraints in (12b)-(12d) are generally non-convex. In particular, the constraint in (12d) represents a Stiefel manifold condition [30], since it enforces orthonormality of the columns of the BD-RIS scattering matrix.
Due to this coupling, a two-stage design is adopted to enhance the sum-rate. In this article, the first stage of the optimization problem is addressed, namely, the design of the scattering matrix. The second stage, concerning the optimization of the transmit beamforming matrices, becomes particularly relevant in more general MU-MIMO settings where active beamforming plays a larger role in the achievable performance. Further discussion on this aspect, as well as on the extension to the general MU-MIMO case, is provided in the next subsection.
II-C A Note on the General MU-MIMO Case
Before we proceed, let us clarify that the method proposed in this article can be directly applied to MU-MIMO systems with relatively few modifications described below.
Consider a MU-cell-free massive MIMO (CF-mMIMO) system with $L$ access points (APs), each equipped with $N_{a}$ TX antennas, such that the total number of transmit antennas is $N_{t}=L\times N_{a}$, and $K$ users equipped with with $M$ RX antennas, aided by a BD-RIS with $R$ REs. The corresponding channel linking the $l$-th AP to the $k$-th user can be defined as
$\mathbf{E}_{l,k}=\mathbf{H}_{\mathrm{RX},k}\mathbf{\Theta}\mathbf{H}_{\mathrm{TX},l}\in\mathbb{C}^{M\times N_{a}},$ (13)
where $\mathbf{H}_{\mathrm{RX},k}$ and $\mathbf{H}_{\mathrm{TX},l}$ denote the BD-RIS-to-user and AP-to-BD-RIS channel matrices, respectively.
Defining the equivalent channel to the $k$-th user as the concatenation of the channel matrices from all APs to the $k$-th user, we write $\mathbf{E}_{k}\triangleq[\mathbf{E}_{1,k},\mathbf{E}_{2,k},\dots,\mathbf{E}_{L,k}]\in\mathbb{C}^{M\times N_{t}}$. Accordingly, the complex baseband received signal at the $k$-th user, which generalizes (1), is given by
$\mathbf{r}_{k}=\overbrace{\,\mathbf{E}_{k}\mathbf{V}_{k}\mathbf{s}_{k}}^{\text{Intended signal}}\;+\!\!\!\!\!\overbrace{\!\!\!\!\!\sum_{k^{\prime}\in\mathcal{K}\backslash\{k\}}\!\!\!\!\!\!\mathbf{E}_{k}\mathbf{V}_{k^{\prime}}\mathbf{s}_{k^{\prime}}}^{\text{Downlink inter-user interference}}\!\!+\mathbf{n}_{k}\in\mathbb{C}^{M\times 1},$ (14)
where $\mathbf{n}_{k}\sim\mathcal{CN}(0,N_{0}\mathbf{I})$.
For a given rate $\log_{2}\text{det}(\mathbf{I}_{M}+\mathbf{\Gamma}_{k})$, the SINR at the $k$-th user, $\mathbf{\Gamma}_{k}$, generalizes (11) and is expressed as
$\mathbf{\Gamma}_{k}=\mathbf{V}_{k}^{\mathrm{H}}\mathbf{E}_{k}^{\mathrm{H}}\mathbf{\Psi}_{k}^{-1}\mathbf{E}_{k}\mathbf{V}_{k}\in\mathbb{C}^{M\times M},$ (15a)
where
[…]
is the interference-plus-noise covariance matrix, and $\mathbf{V}_{k}$ denotes the beamforming matrix for the $k$-th user.
From the above, it is evident that the proposed framework applies directly to general MIMO scenarios, as previously mentioned. We emphasize, however, such a setting would obscure the isolated impact of the BD-RIS scattering matrix design. In addition, a comprehensive treatment of the MIMO case would not be complete without considering also the design of corresponding beamformers. For these reasons, we leave the joint design of TX beamforming and BD-RIS scattering matrix for future work, and continue hereafter addressing the MU-MISO case, for simplicity, but without loss of generality.
III Stage 1: Scattering Matrix Design
We will solve (P1) in an iterative manner. To first design the scattering matrix, the impact of the BS beamforming is fixed by initializing it with a minimum mean square error (MMSE)-based beamformer [^9] Perfect CSI is assumed as discussed previously. However, MMSE precoding has been shown to be robust under imperfect CSI [40], so this assumption does not represent a fundamental weakness of the considered framework.. Accordingly, (2f) is rewritten as
$\mathbf{r}=\mathbf{H}_{\mathrm{RX}}\mathbf{\Theta}\mathbf{H}_{\mathrm{TX}}\mathbf{V}\mathbf{s}+\mathbf{n}\triangleq\mathbf{H}_{\mathrm{RX}}\mathbf{\Omega}\mathbf{V}\mathbf{s}+\mathbf{n}\triangleq\mathbf{E}\mathbf{V}\mathbf{s}+\mathbf{n},$ (16)
where $\mathbf{\Omega}=\mathbf{\Theta}\mathbf{H}_{\mathrm{TX}}=[\boldsymbol{\omega}_{1},\dots,\boldsymbol{\omega}_{N}]\in\mathbb{C}^{R\times N}$ and $\mathbf{E}=\mathbf{H}_{\mathrm{RX}}\mathbf{\Omega}=[\mathbf{e}_{1},\dots,\mathbf{e}_{K}]^{T}\in\mathbb{C}^{K\times N}$ denotes the equivalent channel.
As a consequence, the SINR at the $k$-th user during the design of the scattering matrix is given as
$\hskip-4.30554pt\gamma_{k}=\frac{\left|\mathbf{h}_{k}^{T}\boldsymbol{\Omega}\mathbf{v}_{k}\right|^{2}}{\sum\limits_{i\neq k}\left|\mathbf{h}_{k}^{T}\boldsymbol{\Omega}\mathbf{v}_{i}\right|^{2}+N_{0}}=\frac{\left|\mathbf{e}_{k}\mathbf{v}_{k}\right|^{2}}{\sum\limits_{i\neq k}\left|\mathbf{e}_{k}\mathbf{v}_{i}\right|^{2}+N_{0}}.$ (17)
Making use of (17), the optimization problem of concern can be formulated as
$(\text{P2}):\quad\underset{\mathbf{\Theta}}{\mathrm{maximize}}\quad$ $\sum_{k}\log_{2}(1+\gamma_{k})$ (18a)
subject to $\mathbf{\Theta}\in\mathcal{S}_{a_{1}},$ (18b)
$\mathbf{\Theta}\in\mathcal{S}_{a_{2}}.$ (18c)
III-A Solution for the Group-Connected Architecture
As the group-connected BD-RIS architecture is the most general version considered, it can provide a generic solution to (P2), which in turn can be rewritten as
$(\text{P2a}):\quad\underset{\mathbf{\Theta}}{\mathrm{maximize}}\quad$ $\sum_{k}\log_{2}(1+\gamma_{k})$ (19a)
subject to $\mathbf{\Theta}_{g}=\mathbf{\Theta}_{g}^{T},$ (19b)
$\mathbf{\Theta}_{g}\mathbf{\Theta}_{g}^{H}=\mathbf{I}_{R_{G}},$ (19c)
where $g=\{1,2,\dots,G\}$.
For simplicity, the sum-rate over all users is denoted as
$\eta=\sum_{k}\eta_{k},\quad\text{with}\quad\eta_{k}=\log_{2}(1+\gamma_{k}).$ (20)
As previously noted, the optimization problem (P2a) is inherently non-convex due to the SINR being a ratio of quadratic terms of $\mathbf{\Theta}$. The unitary equality constraint further complicates the problem due to its non-convex nature.
To address this matter, manifold optimization is employed by treating $\mathbf{\Theta}_{g}$ as a point on the Stiefel manifold. Meanwhile, the symmetry constraint restricts $\mathbf{\Theta}_{g}$ to the manifold of symmetric complex matrices, which forms a linear subspace and thus also a smooth manifold. To simultaneously enforce both the unitary and symmetry constraints, an intersection of the two sets is needed. However, as this intersection is challenging to handle directly, a workaround approach is considered by modifying the objective function to add a penalty that promotes symmetry.
Accordingly, the optimization problem then becomes
$\hskip-5.8125pt(\text{P3}):\quad\underset{\mathbf{\Theta}}{\mathrm{maximize}}\quad$ $\!\!\sum_{k}\log_{2}(1+\gamma_{k})-\nu\bigl\|\boldsymbol{\Theta}-\boldsymbol{\Theta}^{T}\bigr\|_{F}^{2}$ (21a)
subject to $\!\!\mathbf{\Theta}_{g}\mathbf{\Theta}_{g}^{H}=\mathbf{I}_{R_{G}},$ (21b)
where $\nu\in\mathbb{R}$ denotes a nonnegative weight [^10] The penalty parameter $\nu$ is selected empirically based on preliminary simulations. While systematic hyperparameter optimization techniques could be employed (see, e.g., [41, 42, 43]), the focus of this work is on the scattering matrix design framework itself rather than on optimal regularization tuning. Moreover, in the following sections, the proposed method is shown to outperform existing SotA approaches even with empirically chosen $\nu$, and further performance improvements may be achieved through adaptive or data-driven tuning strategies. and $\bigl\|\boldsymbol{\Theta}-\boldsymbol{\Theta}^{T}\bigr\|_{F}^{2}$ is used to obtain a quantitative measure of how far the matrix is from being symmetric.
Having defined the optimization problem (P3), its solution can be obtained numerically using the Manopt toolbox [44]. However, measures can be taken to reduce the complexity of the optimization problem. In particular, to address the challenges arising in part from the non-convexity of the sum-rate function, the objective can be transformed into a convex form using FP techniques [45, 46].
To this extent, we apply the Lagrangian Dual Transform (LDT) to the sum-rate objective in (P3), which leads to the equivalent formulation
$\bar{\eta}_{k}=\log_{2}(1+\tau_{k})-\frac{\tau_{k}}{\ln(2)}+\frac{1+\tau_{k}}{\ln(2)}\cdot\frac{\left|\mathbf{e}_{k}\mathbf{v}_{k}\right|^{2}}{\sum\limits_{i}\left|\mathbf{e}_{k}\mathbf{v}_{i}\right|^{2}+N_{0}},\vskip-4.30554pt$ (22)
where the auxiliary variable $\tau_{k}\in\mathbb{C}$ denotes the Lagrange multiplier, s.t. $\tau_{k}=\gamma_{k}$.
The fractional term in the majorized rate expression $\bar{\eta}_{k}$ given in (22) can be further convexified by applying the Quadratic Transform (QT), resulting in the following convex reformulation of the sum-rate objective:
$\hat{\eta}_{k}=$ $\log_{2}(1+\tau_{k})-\frac{\tau_{k}}{\ln(2)}$ (23)
$+\frac{1+\tau_{k}}{\ln(2)}\Big[2\Re\{y_{k}^{\star}\mathbf{e}_{k}\mathbf{v}_{k}\}-\left|y_{k}\right|^{2}\big(\sum_{i}\left|\mathbf{e}_{k}\mathbf{v}_{i}\right|^{2}+N_{0}\big)\Big],$
where $y_{k}\in\mathbb{C}$ is an auxiliary variable introduced by the QT, given as
$y_{k}=\frac{\mathbf{e}_{k}\mathbf{v}_{k}}{\sum\limits_{i}\left|\mathbf{e}_{k}\mathbf{v}_{i}\right|^{2}+N_{0}}.$ (24)
In this manner, the optimization problem (P3) can be reformulated as
$\hskip-5.8125pt(\text{P3a}):\quad\underset{\mathbf{\Theta}}{\mathrm{maximize}}\quad$ $\!\!\sum_{k}\hat{\eta}_{k}-\nu\bigl\|\boldsymbol{\Theta}-\boldsymbol{\Theta}^{T}\bigr\|_{F}^{2}$ (25a)
subject to $\!\!\mathbf{\Theta}_{g}\mathbf{\Theta}_{g}^{H}=\mathbf{I}_{R_{G}},$ (25b)
where, for convenience, the modified sum-rate objective function in (25a), including the penalty term, is denoted as $\breve{\eta}_{k}$.
In addition, to alleviate the high computational burden associated with Manopt, a tailored CGA algorithm [47] is derived in the sequel. This algorithm employs a closed-form expression for the gradient of the objective function, as given in (25a), thereby enabling a more efficient implementation of the scattering matrix design.
To derive a solution for the general group-connected case, we define the vector $\mathbf{h}_{k}^{(g)T}=\mathbf{H}_{RX\left[k,R_{G}(g-1)+1:gR_{G}\right]}$ and the matrix ${\mathbf{H}^{(g)}_{\mathrm{TX}}}=\mathbf{H}_{\mathrm{TX}\left[R_{G}(g-1)+1:gR_{G},1:N\right]}$, which represent the channel components associated with the $g$-th group. Specifically, $\mathbf{h}_{k}^{(g)T}\in\mathbb{C}^{1\times R_{G}}$ denotes the BD-RIS-to-user $k$ channel vector, and ${\mathbf{H}^{(g)}_{\mathrm{TX}}}\in\mathbb{C}^{R_{G}\times N}$ denotes the BS-to-BD-RIS channel matrix corresponding to the group size $R_{G}$, respectively. Within the $g$-th group dimension, the scattering matrix $\mathbf{\Theta}_{g}\in\mathbb{C}^{R_{G}\times R_{G}}$ models the group-specific scattering behavior. Accordingly, the group-wise equivalent channel is defined as $\mathbf{e}_{k}^{(g)}=\mathbf{h}_{k}^{(g)T}\mathbf{\Theta}_{g}{\mathbf{H}^{(g)}_{\mathrm{TX}}}$, where $\mathbf{e}_{k}^{(g)}\in\mathbb{C}^{1\times N}$.
As such, the gradient of the objective function (25a) with respect to each group $\mathbf{\Theta}_{g}$ is computed as
$\nabla_{\mathbf{\Theta}_{g}}\!\breve{\eta}=\nabla_{\mathbf{\Theta}_{g}}\!\Big(\!\sum_{k}\hat{\eta}_{k}-\nu\left\|\mathbf{\Theta}_{g}-\mathbf{\Theta}_{g}^{T}\right\|_{F}^{2}\!\Big),\vskip-4.30554pt$ (26)
where the symmetry penalty term can be rewritten as
$\left\|\mathbf{\Theta}_{g}-\mathbf{\Theta}_{g}^{T}\right\|_{F}^{2}$ $=\mathrm{Tr}\{(\mathbf{\Theta}_{g}-\mathbf{\Theta}_{g}^{T})^{H}(\mathbf{\Theta}_{g}-\mathbf{\Theta}_{g}^{T})\}$ (27)
$=\mathrm{Tr}\{(\mathbf{\Theta}_{g}^{H}-\mathbf{\Theta}_{g}^{\star})(\mathbf{\Theta}_{g}-\mathbf{\Theta}_{g}^{T})\}$
$=\mathrm{Tr}\{\mathbf{\Theta}_{g}^{H}\mathbf{\Theta}_{g}-\mathbf{\Theta}_{g}^{H}\mathbf{\Theta}_{g}^{T}-\mathbf{\Theta}_{g}^{\star}\mathbf{\Theta}_{g}+\mathbf{\Theta}_{g}^{\star}\mathbf{\Theta}_{g}^{T}\}.$
The gradient for half of the terms in (27) can be effectively found in [48, eq. (248)], namely
$\nabla_{\mathbf{\Theta}_{g}}\mathrm{Tr}\{\mathbf{\Theta}_{g}^{H}\mathbf{\Theta}_{g}\}=\nabla_{\mathbf{\Theta}_{g}}\mathrm{Tr}\{\mathbf{\Theta}_{g}^{\star}\mathbf{\Theta}_{g}^{T}\}^{T}=2\mathbf{\Theta}_{g}.$ (28)
Conversely,
$\nabla_{\mathbf{\Theta}_{g}}\mathrm{Tr}\{\mathbf{\Theta}_{g}^{H}\mathbf{\Theta}_{g}^{T}\}=\nabla_{\mathbf{\Theta}_{g}}\mathrm{Tr}\{\mathbf{\Theta}_{g}^{\star}\mathbf{\Theta}_{g}\}^{T},$ (29)
where the gradient is obtained as
$\nabla_{\mathbf{\Theta}_{g}}\mathrm{Tr}\{\mathbf{\Theta}_{g}^{\star}\mathbf{\Theta}_{g}\}=2\frac{\partial\mathrm{Tr}\{\mathbf{\Theta}_{g}^{\star}\mathbf{\Theta}_{g}\}}{\partial\mathbf{\Theta}_{g}^{\star}}=2\mathbf{\Theta}_{g}^{T},$ (30)
using the result from Lemma 1.
The reader is referred to Appendix A for further details.
Based on the above, the gradient of $\left\|\mathbf{\Theta}_{g}-\mathbf{\Theta}_{g}^{T}\right\|_{F}^{2}$ with respect to $\mathbf{\Theta}_{g}$ is derived as
$\nabla_{\mathbf{\Theta}_{g}}\left\|\mathbf{\Theta}_{g}-\mathbf{\Theta}_{g}^{T}\right\|_{F}^{2}=4\left(\mathbf{\Theta}_{g}-\mathbf{\Theta}_{g}^{T}\right).$ (31)
Omitting the constant terms of the equivalent sum-rate objective after the FP transformations allows for the gradient to be expressed as
$\nabla_{\mathbf{\Theta}_{g}}\hat{\eta}_{k}\!=\!\nabla_{\mathbf{\Theta}_{g}}\!\Big(2\Re\{y_{k}^{\star}\mathbf{e}_{k}\mathbf{v}_{k}\}\!-\!\left|y_{k}\right|^{2}\!\big(\sum_{i}\left|\mathbf{e}_{k}\mathbf{v}_{i}\right|^{2}\!+\!N_{0}\big)\Big),\vskip-4.30554pt$ (32)
where the gradient with respect to both terms is given as
$\nabla_{\mathbf{\Theta}_{g}}2\Re\{y_{k}^{\star}\mathbf{e}_{k}\mathbf{v}_{k}\}=2\left(y_{k}^{\star}\frac{\partial\mathbf{e}_{k}\mathbf{v}_{k}}{\partial\mathbf{\Theta}_{g}}\right)^{\star},$ (33)
and
$\nabla_{\mathbf{\Theta}_{g}}\left|y_{k}\right|^{2}\big(\sum_{i}\left|\mathbf{e}_{k}\mathbf{v}_{i}\right|^{2}+N_{0}\big)=2\left|y_{k}\right|^{2}\sum_{i}\left(\big(\mathbf{e}_{k}\mathbf{v}_{i}\big)^{\star}\frac{\partial\mathbf{e}_{k}\mathbf{v}_{i}}{\partial\mathbf{\Theta}_{g}}\right)^{\star},$ (34)
using the result obtained in Lemma 2. The reader is referred to Appendix B for further details.
Finally, the partial derivatives in (33) and (34) are given by
$\hskip-5.38193pt\frac{\partial\mathbf{e}_{k}\mathbf{v}_{l}}{\partial\mathbf{\Theta}_{g}}$ $=\frac{\partial}{\partial\mathbf{\Theta}_{g}}\!\left(\mathbf{h}_{k}^{(g)T}\mathbf{\Theta}_{g}\mathbf{W}^{(g)}\mathbf{v}_{l}\right)\!=\mathbf{h}_{k}^{(g)}(\mathbf{W}^{(g)}\mathbf{v}_{l})^{T}\!,$ (35)
where $l\in\{i,k\}$.
Accordingly, the gradient of the objective function in (26) is reformulated as shown in (36). For convenience, (37) presents the gradient of the objective function assuming a simple power allocation matrix used for beamforming [^11] This assumption can only be made for fully-loaded systems ($K$ = $N$)., $i.e.$, $\mathbf{V}=\mathrm{diag}(\sqrt{v}_{1},\dots,\sqrt{v}_{K})$, where $v_{k}\in\mathbb{R},\forall k$. For convenience, both gradients are provided at the top of the following page.
${\nabla_{\mathbf{\Theta}_{g}}\breve{\eta}=\sum_{k}\frac{1+\tau_{k}}{\ln(2)}\Big[2\left(y_{k}^{\star}\mathbf{h}_{k}^{(g)}\big(\mathbf{W}^{(g)}\mathbf{v}_{k}\big)^{T}\right)^{\star}-2\left|y_{k}\right|^{2}\sum_{i}\left(\big(\mathbf{e}_{k}\mathbf{v}_{i}\big)^{\star}\mathbf{h}_{k}^{(g)}\big(\mathbf{W}^{(g)}\mathbf{v}_{i}\big)^{T}\right)^{\star}\Big]\!-\!4\left(\mathbf{\Theta}_{g}-\mathbf{\Theta}_{g}^{T}\right).}$ (36)
${\nabla_{\mathbf{\Theta}_{g}}\breve{\eta}=\sum_{k}\frac{1+\tau_{k}}{\ln(2)}\Big[2v_{k}\big(y_{k}^{\star}\mathbf{h}_{k}^{(g)}\mathbf{w}_{k}^{(g)T}\big)^{\star}-2\left|y_{k}\right|^{2}\sum_{i}v_{i}^{2}\big(e_{k,i}^{\star}\mathbf{h}_{k}^{(g)}\mathbf{w}_{i}^{(g)T}\big)^{\star}\Big]-4\nu(\mathbf{\Theta}_{g}\!-\mathbf{\Theta}_{g}^{T}).}$ (37)
The constraint of the optimization problem, given in (25b), defines a complex Stiefel manifold of dimension $R_{g}^{2}$. For clarity, a manifold is a topological space that locally resembles the Euclidean space, but may have a globally curved or more complex structure. Such manifolds naturally enforce constraints and symmetries, including orthogonality or unitary conditions [18].
The complex Stiefel manifold corresponding to square blocks ($p=n=R_{G}$) is given by
$\mathrm{St}(R_{G},R_{G};\mathbb{C})=\{\mathbf{\Theta}_{g}\in\mathbb{C}^{R_{G}\times R_{G}}:\mathbf{\Theta}_{g}\mathbf{\Theta}_{g}^{H}=\mathbf{I}_{R_{G}}\}.\vskip-2.15277pt$ (38)
Since the general group-connected BD-RIS scattering matrix is block-diagonal, it can be expressed as
$\hskip-1.89444pt\boldsymbol{\Theta}=\mathrm{blkdiag}(\boldsymbol{\Theta}_{1},\boldsymbol{\Theta}_{2},\dots,\boldsymbol{\Theta}_{G}),\>\boldsymbol{\Theta}_{g}\in\mathrm{St}(R_{G},R_{G};\mathbb{C}),\vskip-2.15277pt$ (39)
where $g\in\{1,2,\dots,G\}$.
The usage of the CGA algorithm is motivated by the rich geometric structure of Riemannian manifolds, which allows well-defined gradients of cost functions. Additionally, optimization over a manifold is locally similar to that in the Euclidean space. As a result, optimization techniques developed for the Euclidean spaces, such as the gradient descent and trust-region methods, have corresponding formulations on manifolds [12].
The algorithm proceeds through a sequence of retraction operations $\mathrm{R}(\cdot,\cdot)$, followed by a projection $\mathrm{T}(\cdot,\cdot)$ onto the tangent space. The retraction operation is expressed as
$\mathrm{R}_{\mathbf{\Theta}}(\mathbf{\Theta},\boldsymbol{\Xi})=\mathcal{Q}(\mathbf{\Theta}+\alpha\boldsymbol{\Xi}),\vskip-2.15277pt$ (40)
where $\alpha$ denotes the step size, $\mathbf{\Theta}$ the current point on the manifold, and $\boldsymbol{\Xi}$ the ascent direction in the tangent space.
In the above, the operator $\mathcal{Q}(\cdot)$ returns the Q-factor from the QR decomposition of $\mathbf{\Theta}+\alpha\boldsymbol{\Xi}$, ensuring that $\mathrm{R}_{\mathbf{\Theta}}(\mathbf{\Theta},\boldsymbol{\Xi})\in\mathrm{St}(R,R;\mathbb{C})$ [49]. Thus, the retraction function maps a point in the tangent space back onto the manifold.
The retraction function for each block is defined as
$\mathbf{Y}_{g}=\mathbf{\Theta}_{g}+\alpha\boldsymbol{\Xi}_{g},\>[\mathcal{Q}_{g},\mathcal{R}_{g}]=\mathcal{QR}(\mathbf{Y}_{g}),\>\mathbf{\Theta}_{g}^{\text{new}}=\mathcal{Q}_{g}.\vskip-2.15277pt$ (41)
Complementary to the retraction operation, the tangent projection function describes the feasible directions of movement at a point on a manifold, represented by tangent vectors. All tangent vectors at a point form the tangent space, which includes every possible direction the point can move to.
Furthermore, the tangent space can be regarded as a Euclidean space, where a particular tangent vector, denoted as the Riemannian gradient, indicates the direction of steepest descent of the objection function [30].
In light of the above, it is initially required for the Euclidean gradient of the objective function to be computed. The Riemannian gradient is then obtained by projecting this Euclidean gradient, as given in (36) and (37), onto the tangent space. This projection is expressed as
$\mathrm{T}_{\mathbf{\Theta}}(\nabla_{\mathbf{\Theta}}\eta,\mathbf{\Theta})=\nabla_{\mathbf{\Theta}}\eta-\mathbf{\Theta}\cdot\frac{\mathbf{\Theta}^{H}\nabla_{\mathbf{\Theta}}\eta+(\nabla_{\mathbf{\Theta}}\eta)^{H}\mathbf{\Theta}}{2}.\vskip-2.15277pt$ (42)
Straightforwardly, for each group $g$, the projection onto the tangent space is given as
$\mathrm{T}_{\mathbf{\Theta}_{g}}(\nabla_{\!\mathbf{\Theta}_{g}}\eta,\mathbf{\Theta}_{g})\!=\!\nabla_{\!\mathbf{\Theta}_{g}}\eta-\mathbf{\Theta}_{g}\cdot\frac{\mathbf{\Theta}_{g}^{H}\nabla_{\mathbf{\Theta}_{g}}\eta+(\nabla_{\mathbf{\Theta}_{g}}\eta)^{H}\mathbf{\Theta}_{g}}{2}.\vskip-2.15277pt$ (43)
After the optimization, $i.e.$, following the sequence of retraction and tangent projection operations, the method from [22] is applied to enforce the symmetry and unitary constraints, since adding a penalty serves only as a workaround approach. The full procedure implemented in this article is outlined in Algorithm 1.
Algorithm 1 Proposed CGA for RBD-RIS Optimization
[…]
0: Optimized RIS matrix $\mathbf{\Theta}^{\mathrm{opt}}$
[…]
1: Compute initial FP auxiliary variables $\tau_{k}$ and $y_{k}$, $\forall k$
2: Compute initial objective ${\breve{\eta}}(\mathbf{\Theta}^{(0)})$ and Riemannian gradient $\mathbf{r}^{(0)}=\mathrm{T}_{\mathbf{\Theta}}(\nabla_{\mathbf{\Theta}}{\breve{\eta}}(\mathbf{\Theta}^{(0)}),\mathbf{\Theta}^{(0)})$ using (42)
3: Set initial search direction $\boldsymbol{\Xi}^{(0)}=-\mathbf{r}^{(0)}$
4:for $i=0$ to II do
5:  if $\langle\mathbf{r}^{(i)},\boldsymbol{\Xi}^{(i)}\rangle\leq 0$ then
6:   Set $\boldsymbol{\Xi}^{(i)}=\mathbf{r}^{(i)}$
7:  end if
8:  Compute $\alpha^{(i)}$ via Armijo line search and update $\mathbf{\Theta}^{(i+1)}$
9:  Compute new FP auxiliary variables $\tau_{k}$ and $y_{k}$, $\forall k$
10:  Compute new objective ${\breve{\eta}}(\mathbf{\Theta}^{(i+1)})$, using (25a)
11:  Compute new Riemannian gradient $\mathbf{r}^{(i+1)}$, using (43)
12:  Compute $\beta^{(i)}=\max\left(0,\frac{\langle\mathbf{r}^{(i+1)},\mathbf{r}^{(i+1)}-\mathbf{r}^{(i)}\rangle}{\langle\mathbf{r}^{(i)},\boldsymbol{\Xi}^{(i)}\rangle}\right)$
13:  Update direction: $\boldsymbol{\Xi}^{(i+1)}=-\mathbf{r}^{(i+1)}+\beta^{(i)}\boldsymbol{\Xi}^{(i)}$
14:  if $|\eta(\mathbf{\Theta}^{(i+1)})-\eta(\mathbf{\Theta}^{(i)})|<\epsilon$ then
15:   break
16:  end if
17:end for
18:for $g=1$ to $G$ do
19:  Extract block $\mathbf{\Theta}_{g}$ from $\mathbf{\Theta}^{(i+1)}$
20:  Symmetrize: $\mathbf{\Theta}_{\mathrm{sym}}=\frac{1}{2}(\mathbf{\Theta}_{g}+\mathbf{\Theta}_{g}^{T})$
21:  Perform SVD: $\mathbf{\Theta}_{\mathrm{sym}}=\mathbf{U}\mathbf{\Sigma}\mathbf{V}^{H}$
22:  Set $\mathbf{\Theta}^{\mathrm{opt}}_{g}=\mathbf{U}\mathbf{V}^{H}$
23:end for
24: Assemble full $\mathbf{\Theta}^{\mathrm{opt}}$ as block diagonal of all $\mathbf{\Theta}_{g}^{\mathrm{opt}}$
25:return $\mathbf{\Theta}_{\mathrm{opt}}$
The algorithm begins by initializing a random unitary symmetric block-diagonal matrix $\mathbf{\Theta}^{(0)}$ and computing the initial objective function along with its gradient. In each iteration, these quantities are updated according to the current values of the scattering matrix until convergence is achieved. The step size $\alpha$ is determined using an Armijo line search [30, 50], which ensures that there is a sufficient ascent. Furthermore, the parameter $\beta$ denotes the conjugate gradient (CG) momentum term, used to enhance the convergence speed of the algorithm.
Among various options, including the Fletcher-Reeves, Polak Ribière, and Hestenes-Stiefel formula, this algorithm incorporates the Polak-Ribière formula as described in [47, 50]. Finally, the scattering matrix resulting from the series of retraction and tangential projection operations is further processed using the method described in [22], ensuring that $\mathbf{\Theta}_{g}^{\mathrm{opt}}$ satisfies both the symmetric and unitary constraints.
Computational Complexity: The computational complexity of Algorithm 1 is primarily influenced by the Armijo line search. Within the line search, two operations dominate the complexity, namely a triple matrix multiplication and a QR decomposition-based retraction. Considering the general case of the group-connected architecture, the QR decomposition for each group, of size $R/G$, has a complexity of $\mathcal{O}\big((R/G)^{3}\big)$, while the triple matrix multiplication has a complexity of $\mathcal{O}\big(KR(R+N)\big)$. Since the QR decomposition is applied independently to each group, the total complexity of the retraction operation across all groups and users is given as $\mathcal{O}\big(KG(R^{3}/G^{3})\big)$.
Table I: Computational Complexity of BD-RIS Scattering Matrix Design
BD-RIS Architecture Computational Complexity
Interference Nulling SotA Alg. [21] Proposed Alg. 1
Single-connected $\mathcal{O}(tR+K^{6}+K^{4}R+K^{2}R^{2}+K^{2}R)$ $\mathcal{O}(IL(KR+K^{2}R^{2}+K^{2}RN))$
Group-connected $\mathcal{O}(tG\!R_{G}^{3}\!+\!tG\!R_{G}^{2}\!+\!K^{6}\!\!+\!K^{4}G\!R_{G}^{2}\!+\!K^{2}\!G^{2}\!R_{G}^{4}\!+\!K^{2}\!G\!R_{G}^{2})$ $\mathcal{O}(IL(KGR_{G}^{3}+K^{2}R^{2}+K^{2}RN))$
Fully-connected $\mathcal{O}(tR^{3}+tR^{2}+K^{6}+K^{4}R^{2}+K^{2}R^{4}+K^{2}R^{2})$ $\mathcal{O}(IL(KR^{3}+K^{2}R^{2}+K^{2}RN))$
In contrast, the triple matrix multiplication is performed on the full system, $i.e.$, not on a per-group basis, and yields a total complexity for all users of $\mathcal{O}\big(K(KR^{2}+KRN)\big)$. With $L$ Armijo iterations, these contributions accumulate, giving an overall line search complexity of
$\mathcal{O}\big(LKG(R^{3}/G^{3})+LK(KR^{2}+KRN)\big).\vskip-2.15277pt$ (44)
The gradient computation consists of several terms, resulting in a total complexity across all groups and users of $\mathcal{O}\big(K^{2}G(R/G)^{2}\big)$, and the computation of the FP auxiliary variables, which further amounts to $\mathcal{O}\big(K^{2}N\big)$. Furthermore, the initialization and projection described in [22] have computational complexity $\mathcal{O}\big(R^{3}/G^{2}\big)$, which stems from the singular value decomposition (SVD) computation. Lastly, the tangential projection in (43) has a complexity of $\mathcal{O}\big(R^{3}\big)$. Considering $I$ iterations of the algorithm, and the case $L\gg 1$, the total complexity of the algorithm can then be expressed as
$\mathcal{O}\big(IL(KR^{3}/G^{2}+K^{2}R^{2}+K^{2}RN)\big).\vskip-2.15277pt$ (45)
Table I summarizes the simplified computational complexity associated with the different BD-RIS architectures, enabling a direct comparison with the SotA method from [21] [^12] Notice that the aforementioned SotA technique assumes that the number of TX antennas at the BS equals the number of single-antenna users ($K=N$). The proposed method does not require such assumption and therefore can be considered a generalization of the latter in this regard., where $t$ denotes the number of iterations required by the interference nulling algorithm [21]. In addition, we compare the proposed method against the joint scattering and beamforming matrix design approach based on the pp-ADMM framework [34]. Accounting for $T_{\mathrm{ADMM}}$ iterations required to satisfy the convergence criterion, the overall computational complexity of the pp-ADMM algorithm scales as
$\mathcal{O}\Big(T_{\mathrm{ADMM}}\Big[R^{3}K+\big(\min\!\big\{\sum\nolimits_{i=1}^{R}|S_{i}|,\,2RK\big\}\big)^{3}\Big]\Big),\vskip-2.15277pt$ (46)
where $S_{i}$ denotes the set of non-zero entries ($i.e.$, active connections) in the $i$-th row of the BD-RIS scattering matrix, such that $\sum_{i=1}^{R}|S_{i}|$ represents the total number of independent optimization variables determined by the connectivity structure of the BD-RIS.
Figure 3: Convergence of Algorithm 1 vs. number of iterations II for different connectivity structures, with $P_{\mathrm{max}}=20\,\mathrm{dBm}$, $K=2$, $N=2$, $R=32$, and $K$ factor $K_{f}=$-$\infty$ dB.
The convergence behavior of Algorithm 1 is shown in Figure 3 for the different BD-RIS architectures considered, $i.e.$, single-, group-, with group size of 2 and 4, and fully-connected, respectively. For clarity, the figure represents results generated for one Monte-Carlo iteration, hence the reason for the non-smooth convergence curves.
Each BD-RIS architecture is represented by a distinct marker, as indicated in the legend of Figure 3, and this notation is used consistently throughout the article to facilitate cross-referencing of results.
The simulation settings are as follows: convergence tolerance of $\epsilon=10^{-8}$, maximum number of CGA iterations $I=8000$, maximum Armijo line search steps $L=200$, sufficient increase coefficient for the stepsize of $2\times 10^{-11}$, initial stepsize $c_{\text{init}}=1$, contraction factor for the stepsize $c_{\text{dec}}=0.75$, and the symmetry enforcing penalty constant $\nu=1$. Results indicate that more complex architectures require a larger number of iterations to converge; for instance, the single-connected case converges in approximately $100$ iterations, while the full-connected case requires around $600$ iterations under the same conditions.
Additionally, the convergence speed of the algorithm depends on the number of REs and the maximum transmit power $P_{\mathrm{max}}$, as well as on the initial scattering matrix, which in this article is chosen as a random unitary symmetric matrix. Regarding solution quality, since the reformulated optimization problem in (25a) is convex, the FP-based approach applied to the originally non-convex problem is guaranteed to converge to a stationary point, as the adopted approximation yields a monotonically non-decreasing objective sequence [45, 46]. However, due to the non-convexity of the original problem, the stationary point generally corresponds to a local optimum.
As validated in Figure 3, the number $I$ of iterations required for convergence is considered for the scenario where $P_{\mathrm{max}}=20\mathrm{dBm}$, $K=2$ users, $N=2$ TX antennas, and $R=32$ REs, under Rayleigh fading conditions, $i.e.$, with a $K$-factor $(K_{f})$ corresponding to -$\infty$ dB. Specifically, these are $I_{\mathrm{SC}}=100$, $I_{\mathrm{GC}_{2}}=200$, $I_{\mathrm{GC}_{4}}=400$ and $I_{\mathrm{FC}}=600$, corresponding to the single-, group-, with group size 2 and 4, and fully-connected architectures, respectively.
As can be learned from Table I and (46), our proposed algorithm offers scalability, since its computational complexity grows with lower exponents on key system parameters such as the number of users $K$, the number of RIS elements $R$ and the number of BS antennas $N$. This advantage is in addition to the superior communications performance, which shall be thoroughly demonstrated in the next section.
For systems of smaller sizes, however, the performance advantage of the proposed method may come at the cost of a higher complexity compared to the SotA techniques [21, 34], due to the large number of iterations required for convergence of the algorithm, as shown in Figure 3. The improved sum-rate and increased robustness observed across multiple BD-RIS architectures, resulting from the algorithm’s design, are thoroughly presented and analyzed in the subsequent section.
IV Simulation Results
This section aims to showcase the effectiveness of the proposed BD-RIS-aided system in enhancing communication performance via computer simulations executed with the following parameters. Unless otherwise specified, it will be assumed throughout the simulations that the BS is equipped with $N=2$ TX antennas and serves $K=2$ single RX antenna users [^13] This choice is motivated by the goal of isolating and evaluating the performance gains introduced by the BD-RIS. In more heavily loaded scenarios, these gains may be partially masked by additional beamforming gains., the RIS structures consist of $R=32$ REs and the maximum transmit power $P_{\mathrm{max}}$ ranges from 0 to $20\,\mathrm{dBm}$. Furthermore, the parameters utilized in Algorithm 1 are the same as those previously used to generate the results shown in Figure 3. In addition to evaluating the proposed scattering matrix design, the numerical results present the achieved sum-rate when the optimized BD-RIS is combined with a fixed MMSE-based beamformer. Joint optimization of the scattering matrix and beamforming is not considered, allowing the impact of the proposed scattering matrix design to be assessed independently.
(a) Single-connected
(b) Group-connected
(c) Fully-connected
Figure 4: Comparison of sum-rate performance of the proposed vs. SotA [21, 34] scattering matrix design with uniform power allocation, considering the single-connected “SC”, group-connected “GC”, with group sizes of 2 and 4, “GC(2)” and “GC(4)”, and the fully-connected “FC” architecture.
The channel model adopted in this article considers both i.i.d. Rayleigh and Rician small-scale fading with distance-dependent large-scale attenuation expressed as $\Upsilon(d)=C_{0}(\frac{d}{d_{0}})^{-\rho}$, where $C_{0}=-30\,\mathrm{dB}$ denotes the reference channel gain at a reference distance $d_{0}=1$ m, $\rho=2.2$ is the path-loss exponent, $d=50$ m is the distance between the BS and BD-RIS, and $d=2.5$ m is the distance between the BD-RIS and the users [^14] Although i.i.d fading channels are not entirely consistent with two-dimensional (2D) surfaces, as shown in [51], this model is adopted here only to enable a fair comparison with the SotA methods [21, 34], since the scattering matrix design itself does not require any specific model to be feasible. The fading parameters and large-scale attenuation settings are selected consistently with the aforementioned SotA methods to ensure meaningful performance comparisons. Importantly, the proposed scattering matrix design framework does not rely on any specific small-scale fading assumption.. Specifically, some numerical results assume Rician fading with $K_{f}=2\,\mathrm{dB}$, while others consider the Rayleigh case corresponding to $K_{f}=-\infty\,\mathrm{dB}$. Lastly, the system is assumed to operate at 2.42.4GHz, and each user experiences identical noise with a noise power $N_{0}=-80\mathrm{dBm}$.
The first set of results is presented in Figure 4 to demonstrate the effectiveness of the proposed method in enhancing system performance. Different architectures, such as the single-, group-, and fully-connected are considered, and results comparing the sum-rate performance achieved by the proposed and the SotAtechniques [21, 34] are provided. Notably, the results indicate that the proposed scattering matrix “SM”, combined with either uniform power allocation “PA”, or a fixed MMSE-based beamformer “BF”, outperforms both SotAcounterparts considered.
In particular, the interference nulling SM design with uniform PA [21] and the joint scattering and beamforming matrix design based on the pp-ADMM framework [34], are outperformed with all group architectures, under Rayleigh fading conditions with $K_{f}=$-$\infty$ dB and for a small number of users and TX antennas at the BS, i.e., $K=2$ and $N=2$.
Figure 5: CDF of sum-rate performance of the proposed vs. SotA [21, 34] scattering matrix design with uniform power allocation, considering the fully-connected “FC”, group-connected “GC”, with group sizes of 2 and 4, “GC(2)” and “GC(4)”, and the single-connected ”SC” architecture.
(a) Single-connected
(b) Group-connected
(c) Fully-connected
Figure 6: Comparison of sum-rate performance using different beamforming schemes, considering the proposed and SotA [21, 34] single-connected “SC”, group-connected “GC”, with group sizes of 2 and 4, “GC(2)” and “GC(4)”, and the fully-connected “FC” architecture.
Complementing the assessment in terms of sum-rate performance, Figure 5 further shows the cumulative distribution function (CDF) of the sum-rate performance of the proposed BD-RIS scattering matrix design, alongside the SotA methods [21, 34] for all considered architectures, $i.e.$, single-, group- with group sizes of 2 and 4, and fully-connected.
More specifically, the group-connected architecture with group size 4, when using the proposed method, clearly outperforms the considered SotA designs with the same architecture. The fully-connected architecture with the proposed method achieves, on average, slightly better performance than [34]. Architectures with lower complexity, namely the single-connected architecture and the group-connected architecture with group size 2, achieve slightly lower performance than [34]; however, all proposed architectures still outperform the SotA interference nulling design in [21].
Next, Figure 6 presents a comparison of the sum-rate performance of the proposed scattering matrix “SM”, in combination with either uniform power allocation “PA”, or a fixed MMSE-based beamformer “BF”, against the corresponding SotA schemes, namely the interference nulling SM design with uniform PA [21] and the joint scattering and beamforming matrix design based on the pp-ADMM framework [34].
The comparison is conducted across all considered architectures under Ricean fading conditions with $K_{f}=2$ dB, and for a small number of users and TX antennas at the BS, specifically $K=2$ and $N=2$, highlighting the robustness of the proposed method across different fading conditions. We remark that notable performance gains are particularly observed for the group-connected architecture, with group size 2, where the proposed method outperforms both SotA methods, even those with a higher connectivity, by an order of 2, in the high transmit power regime.
Figure 7: Comparison of sum-rate performance of the proposed vs. SotA [21, 34] methods as a function of the number of REs, i.e. $R$, considering the fully-connected “FC”, group-connected “GC”, with group sizes of 2 and 4, “GC(2)” and “GC(4)”, and the single-connected ”SC” architecture.
Finally, Figure 7 presents the sum-rate performance of the proposed method, in combination with uniform power allocation “PA”, in comparison to the SotA methods [21, 34], as a function of the number of REs $R$. This result further builds upon the claim that for all BD-RIS architectures considered, $i.e.$, single-, group-, and fully-connected, and under the Rayleigh fading scenario, the proposed scattering matrix consistently outperforms the SotA benchmarks [21, 34]. Importantly, the performance gain is maintained as the number of reflective elements increases, which highlights the robustness of the proposed approach, and showcases that the proposed solution is scalable.
V Conclusion
A novel sum-rate maximization scheme for RBD-RIS-aided MU-MISO systems is presented. To this end, an optimization problem was formulated to determine the BD-RIS scattering matrix configuration that maximizes the users’ sum-rate, and solved using a Riemannian manifold steepest descent method with a closed-form gradient. Simulation results validated the effectiveness of the proposed approach, showing consistent performance gains over the SotA, where the improvements arise from jointly exploiting the geometric structure of the manifold and the direct sum-rate based problem formulation. Additionally, the simulation results were compared against joint scattering matrix and beamforming designs. Under the considered scenarios, the proposed method outperformed the design with fixed beamforming, which opens up the possibility for beamforming optimization.
This is promising, as further beamforming design optimization is expected to provide even greater performance improvements. Overall, the results show that manifold-based scattering matrix optimization enables scalable and robust performance in RIS-aided systems. Future work will extend the framework to include transmit beamforming design.
Appendix A Proof of Lemma 1
Lemma 1.
Consider $\mathbf{A}\in\mathbb{C}^{N\times N}$ with entries defined as
[…]
Then
$\nabla_{\mathbf{A}}\mathrm{Tr}\{\mathbf{A}^{\star}\mathbf{A}\}=2\frac{\partial\mathrm{Tr}\{\mathbf{A}^{\star}\mathbf{A}\}}{\partial\mathbf{A}^{\star}}=2\mathbf{A}^{T},\vskip-4.30554pt$ (48)
where
$\mathrm{Tr}\{\mathbf{A}^{\star}\mathbf{A}\}=\sum_{i=1}^{N}\sum_{k=1}^{N}a_{i,k}^{\star}a_{k,i},$ (49)
with $a_{i,k}^{\star}$ and $a_{k,i}\in\mathbb{C}$, such that they can be written in the form $x=\Re{x}+j\Im{x}$.
Proof: For simplicity, let $\phi=\mathrm{Tr}\{\mathbf{A}^{\star}\mathbf{A}\}$. Then,
$\frac{\partial\phi}{\partial\Re{\mathbf{A}}}={\begin{pmatrix}\frac{\partial\phi}{\partial\Re a_{1,1}}&\cdots&\frac{\partial\phi}{\partial\Re a_{1,N}}\\[-4.30554pt] \vdots&\ddots&\vdots\\ \frac{\partial\phi}{\partial\Re a_{N,1}}&\cdots&\frac{\partial\phi}{\partial\Re a_{N,N}}\end{pmatrix}=2\Re\mathbf{A}^{T},}\vskip-4.30554pt$ (50)
and
$j\frac{\partial\phi}{\partial\Im{\mathbf{A}}}={\begin{pmatrix}\frac{\partial\phi}{\partial\Im a_{1,1}}&\cdots&\frac{\partial\phi}{\partial\Im a_{1,N}}\\[-4.30554pt] \vdots&\ddots&\vdots\\ \frac{\partial\phi}{\partial\Im a_{N,1}}&\cdots&\frac{\partial\phi}{\partial\Im a_{N,N}}\end{pmatrix}=2j\Im\mathbf{A}^{T}.}\vskip-4.30554pt$ (51)
Straightforwardly, the Generalized Complex Derivative and Conjugate Complex Derivative [48, eq.(229), eq.(230)] are obtained as
$\frac{\mathrm{d}\phi}{\mathrm{d}\mathbf{A}}=\frac{1}{2}\Big(2\Re\mathbf{A}^{T}-2j\Im\mathbf{A}^{T}\Big)=\mathbf{A}^{H},$ (52)
$\frac{\mathrm{d}\phi}{\mathrm{d}\mathbf{A}^{\star}}=\frac{1}{2}\Big(2\Re\mathbf{A}^{T}+2j\Im\mathbf{A}^{T}\Big)=\mathbf{A}^{T}.$ (53)
As evidence that $\phi$ is a real-valued function, the following is provided
$\mathrm{Tr}\{\mathbf{A}^{\star}\mathbf{A}\}=\sum_{i}^{N}|a_{i,i}|^{2}+2\Re\Big\{\sum_{i}^{N}\sum_{k=i+1}^{N}a_{i,k}^{\star}a_{k,i}\Big\},\vskip-4.30554pt$ (54)
as such, the Complex Gradient Matrix [48, eq.(233)] can be computed as follows
$\nabla_{\mathbf{A}}\mathrm{Tr}\{\mathbf{A}^{\star}\mathbf{A}\}=2\frac{\partial\mathrm{Tr}\{\mathbf{A}^{\star}\mathbf{A}\}}{\partial\mathbf{A}^{\star}}=2\mathbf{A}^{T},\vskip-4.30554pt$ (55)
hence, (30) is proven. $\square$
Appendix B Proof of Lemma 2
Lemma 2.
Consider $\mathbf{a}\in\mathbb{C}^{K\times 1}$, $\mathbf{B}\in\mathbb{C}^{K\times K}$, and $\mathbf{c}\in\mathbb{C}^{K\times 1}$, such that $d=\mathbf{a}^{T}\mathbf{B}\mathbf{c}$, where $d\in\mathbb{C}$ and
$d={\left(\!\begin{array}[]{c}a_{1}\\[-4.30554pt] \vdots\\[-4.30554pt] a_{N}\end{array}\!\right)^{T}\!\!\!\left(\!\begin{array}[]{cccc}B_{1,1}&\cdots&B_{1,N}\\[-4.30554pt] \vdots&\ddots&\vdots\\ B_{N,1}&\cdots&B_{N,N}\end{array}\!\right)\!\!\!\left(\!\begin{array}[]{c}c_{1}\\[-4.30554pt] \vdots\\[-4.30554pt] c_{N}\end{array}\!\right).}\vskip-4.30554pt$ (56)
Consider the function $f(d)=|d|^{2}$. The gradient of $f(d)$ with respect to $\mathbf{B}$ is given as
$\nabla_{\mathbf{B}}|d|^{2}=d\mathbf{a}^{\star}\mathbf{c}^{H},\vskip-4.30554pt$ (57)
where $\mathbf{a}=\Re\mathbf{a}+j\Im\mathbf{a}$, $\mathbf{B}=\Re\mathbf{B}+j\Im\mathbf{B}$ and $\mathbf{c}=\Re\mathbf{c}+j\Im\mathbf{c}$, where $\Re\backslash\Im\mathbf{a}$, $\Re\backslash\Im\mathbf{c}\in\mathbb{R}^{K\times 1}$ and $\Re\backslash\Im\mathbf{B}\in\mathbb{R}^{K\times K}$.
Proof.
Following the steps for calculating the gradient of complex valued functions, as described in [48, eq.(229), eq.(230), eq.(233)]. The derivatives with respect to the real and imaginary parts of $\mathbf{B}$ are calculated, as shown below
$\frac{\partial|d|^{2}}{\partial\Re\mathbf{B}}=\frac{\partial|d|^{2}}{\partial d}\frac{\partial d}{\partial\Re\mathbf{B}}+\frac{\partial|d|^{2}}{\partial d^{\star}}\frac{\partial d^{\star}}{\partial\Re\mathbf{B}},\vskip-4.30554pt$ (58)
and
$\frac{\partial|d|^{2}}{\partial\Im\mathbf{B}}=\frac{\partial|d|^{2}}{\partial d}\frac{\partial d}{\partial\Im\mathbf{B}}+\frac{\partial|d|^{2}}{\partial d^{\star}}\frac{\partial d^{\star}}{\partial\Im\mathbf{B}}.\vskip-4.30554pt$ (59)
Straightforwardly
$\frac{\partial|d|^{2}}{\partial d}=d^{\star}\quad\text{and}\quad\frac{\partial|d|^{2}}{\partial d^{\star}}=d.$ (60)
Expanding $d=\mathbf{a}^{T}\mathbf{B}\mathbf{c}$ and $d^{\star}=\mathbf{c}^{H}\mathbf{B}^{H}\mathbf{a}^{\star}$ as a product of complex numbers. The following derivatives are obtained
$\frac{\partial d}{\partial\Re\mathbf{B}}={\begin{pmatrix}\frac{\partial d}{\partial\Re B_{1,1}}&\cdots&\frac{\partial d}{\partial\Re B_{1,N}}\\[-4.30554pt] \vdots&\ddots&\vdots\\ \frac{\partial d}{\partial\Re B_{N,1}}&\cdots&\frac{\partial d}{\partial\Re B_{N,N}}\end{pmatrix}=\mathbf{a}\mathbf{c}^{T},}$ (61)
$\frac{\partial d}{\partial\Im\mathbf{B}}={\begin{pmatrix}\frac{\partial d}{\partial\Im B_{1,1}}&\cdots&\frac{\partial d}{\partial\Im B_{1,N}}\\[-4.30554pt] \vdots&\ddots&\vdots\\ \frac{\partial d}{\partial\Im B_{N,1}}&\cdots&\frac{\partial d}{\partial\Im B_{N,N}}\end{pmatrix}=j\mathbf{a}\mathbf{c}^{T}.}$ (62)
Similarly, for $d^{\star}$
$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{\partial d^{\star}}{\partial\Re\mathbf{B}}=\mathbf{a}^{\star}\mathbf{c}^{H}\quad\text{and}\quad\frac{\partial d^{\star}}{\partial\Im\mathbf{B}}=-j\mathbf{a}^{\star}\mathbf{c}^{H}.$ (63)
Finally we obtain
$\frac{\partial|d|^{2}}{\partial\Re\mathbf{B}}=d^{\star}\mathbf{a}\mathbf{c}^{T}+d\mathbf{a}^{\star}\mathbf{c}^{H},$ (64)
$\frac{\partial|d|^{2}}{\partial\Im\mathbf{B}}=jd^{\star}\mathbf{a}\mathbf{c}^{T}-jd\mathbf{a}^{\star}\mathbf{c}^{H},$ (65)
$\frac{\mathrm{d}|d|^{2}}{\mathrm{d}\mathbf{B}^{\star}}=\frac{1}{2}\Big(\frac{\partial|d|^{2}}{\partial\Re\mathbf{B}}+j\frac{\partial|d|^{2}}{\partial\Im\mathbf{B}}\Big)=d\mathbf{a}^{\star}\mathbf{c}^{H}.$ (66)
As the absolute value squared of a complex number is a real value, $i.e.$, $|d|^{2}$ is a real function, the gradient of $f(d)$ with respect to $\mathbf{B}$ is given as
$\nabla_{\mathbf{B}}|d|^{2}=2\frac{\mathrm{d}|d|^{2}}{\mathrm{d}\mathbf{B}^{\star}}=2d\mathbf{a}^{\star}\mathbf{c}^{H}=2\big(d^{\star}\mathbf{a}\mathbf{c}^{T}\big)^{\star},\vskip-4.30554pt$ (67)
hence, (33) and (34) are proven. ∎